### v1

# von Neumann entropy and quantum version of thermodynamic entropy

Abstract
The debate whether the von Neumann (VN) entropy is a suitable quantum version of the thermodynamic (TH) entropy has a long history. In this regard, we briefly review some of the main reservations about the VN entropy and explain that the objections about its time-invariance and subadditivity properties can be either avoided or addressed convincingly. In a broader context, we analyze here whether and when the VN entropy is the same or equivalent to the TH entropy. For a thermalized isolated or open system, the VN entropy for an appropriately chosen density operator is the same as the quantum version of, respectively, the Boltzmann or Gibbs entropy (these latter entropies are equivalent to the TH entropy for large thermalized systems). Since the quantum thermalization is essentially defined for a subsystem of a much larger system, it is important to investigate if the VN entropy of a subsystem is equivalent to its TH entropy. We here show that the VN entropy of a small subsystem of a large thermalized system is proportional to the quantum statistical entropy (Boltzmann’s version) of the subsystem. For relevant numerical results, we take a one-dimensional spin-1/2 chain with next-nearest neighbor interactions.
preprint: APS/123-QED
I Introduction
Since the development of quantum mechanics, there has been an enormous progress in our understanding of how the quantum many-body systems behave. This progress saw some foundational questions being asked and at-least partially answered. One of such questions is on whether and how the statistical mechanics works for the quantum many-body systems[1, 2, 3]. This gives rise to the important field of quantum statistical mechanics (QSM). The thermodynamic (TH) entropy is a very important state function which helps us understand how a system approaches equilibrium besides playing a crucial role in studying equilibrium properties of a system [4]. The usefulness of the TH entropy naturally calls for its appropriate quantum counterpart in QSM.
The TH entropy, as defined by Clausius in the context of second law of thermodynamics, is a state function which takes the maximum value for a system in the equilibrium under fixed macroscopic conditions [4, 5]. For an isolated system, this entropy never decreases when the system undergoes some process. Thus any entropy defined in the similar spirit for a quantum system must be a state function and should increase (or remain constant) when the system undergoes a process.
Boltzmann, and later Gibbs, showed us how the TH entropy can be calculated from the microscopic details using statistical approaches [4, 6]. For isolated system, Boltzmann gave formula (actually Planck first wrote down the explicit expression for the Boltzmann entropy) for the TH entropy in terms of the volume of the relevant phase space (or the number of associated microstates). For the open system, Gibbs on the other hand, using his ensemble theory, provided the formula for the TH entropy in terms for probability distribution function in phase space. The quantum versions of these statistical entropies (Boltzmann’s and Gibbs’) are somewhat intuitively clear [7, 5]: for an isolated quantum system, the entropy is given in terms of the dimension of the relevant subspace of the system’s Hilbert space while for an open quantum system, the entropy can be calculated from the partition function (which depends on the eigenvalues of the appropriate Hamiltonian of the system). Question, however, remains on the quantum version of the TH entropy for a part (subsystem) of a quantum system. We will address this issue separately in this work.
From the thermodynamical considerations, von Neumann (VN) gave a formula for entropy of a quantum ensemble (representing a given mixed state) [8]. This entropy, known as VN entropy, attracted a lot of attention and criticism in regards to its connection to the TH entropy. We show in this work that, for isolated and open quantum system or for small part of a large quantum system, the VN entropy is equivalent to the TH entropy if the quantum system is large and in thermal equilibrium. We largely exclude from our discussions the quantum systems which do not thermalize or systems not in equilibrium.
It is worth mentioning here the logical steps we follow in this work while establishing the said equivalence between the entropies and finding the conditions under which the equivalence works. First, identify under what conditions the TH entropy is equivalent to the statistical entropy (either Boltzmann’s or Gibbs’ version), then reinterpret the statistical entropy in terms of the quantum variables (in Boltzmann’s case: dimension of relevant subspace in place of volume of accessible phase-space, and in Gibbs’ case: energy eigenvalues in place of continuous classical energy). In the last step of understanding the equivalence, one checks if the VN entropy of an appropriate density matrix is equivalent to the quantum version of the statistical entropy (i.e. the reinterpreted statistical entropy in terms of quantum variables). To better understand the equivalence, one should also check if the the VN entropy inherits the important properties of the equilibrium TH entropy. Accordingly, whereever appropriate, we briefly deliberate, if the VN entropy is a state function, additive and whether it increases for isolated system due to an irreversible process.
We organize our paper in the following way. In the next section (Sec. II), we review the thermodynamic and statistical entropies, and subsequently, discuss the suitable quantum versions of these statistical entropies. In Sec. III, we discuss von Neumann entropy and some reservations about the entropy while it comes to representing the thermodynamic entropy. Next, in Sec. IV, we check for three different cases (isolated system, open system and subsystem) if the von Neumann entropy is equivalent to the thermodynamic entropy. We conclude our work in Sec. V.
II Thermodynamic and statistical entropies
In this section we first briefly review the contexts and definitions of the thermodynamic (TH) entropy and the statistical (ST) entropies (both Boltzmann’s and Gibbs’ versions). Next we try to understand under what conditions these entropies (TH and ST) are equivalent. At the end this section, we discuss the reinterpretations of the statistical entropies in terms of the quantum variables.
II.1 Entropy: from Clausius to Boltzmann and Gibbs
The Clausius theorem or the Clausius inequality states that [4], for any cyclic transformation or thermodynamic process, $\oint\frac{d\bar{}\hskip 0.70004ptq}{T}\leq 0$, where $d\bar{}\hskip 1.00006ptq$ is the heat increment received by the system at temperature $T$. Here the equality sign works only for a reversible cycle and this helps us to define a thermodynamic (TH) state function, called entropy. Streaky speaking, this state function (thermodynamic entropy, $S_{TH}$) is well defined only for a thermodynamic system in equilibrium. If a system in equilibrium at temperature $T$ receives $d\bar{}\hskip 1.00006ptq$ amount of heat reversibly and subsequently reaches an equilibrium again, then the difference in entropy between these two equilibrium phases would be $dS_{TH}=\frac{d\bar{}\hskip 0.70004ptq}{T}$.
The Boltzmann’s version of statistical entropy is defined by the number of accessible microstates $\Omega$ of a system (with a fixed energy). For a gaseous system, this number is proportional to the phase-space volume available to the system in the given macrostate. Here the statistical entropy is defined as $S_{BN}=k_{B}\ln{\Omega}$ ) with $k_{B}$ being the Boltzmann (BN) constant [4, 9, 7]. By dividing the accessible phase-space as per the values of the well-defined macrovariables, one can in-fact calculate the BN entropy for each such division [7, 5]. This way one can even make the BN entropy useful in studying a system approaching equilibrium.
[…]
In the following subsection we discuss under what conditions the TH entropy and the statistical entropy (BN or GB entropy) are equivalent.
II.2 Conditions of equivalence between thermodynamic and statistical entropy
To understand the conditions under which the equivalence works, we outline here a method to establish such equivalence. A common way of showing the equivalence between the TH and BN entropies of a system with a given energy is to bring the system with energy $E_{1}$ in contact with another system with energy $E_{2}$ [6]. Before the contact, the systems were in thermal equilibrium and after the contact they will again come to a common equilibrium via exchange of heat. Although the systems exchange heat, total energy ($E=E_{1}+E_{2}$) of the composite system remains constant; here one ignores the interaction energy between the systems. In the physical setup to establish the equivalence, the main premise is that a thermodynamic system in equilibrium with a given energy can be in any of a certain number of microstates with an equal probability (the principle of equal a priori probabilities). Here the number of accessible microstates is a function of the system’s energy. If $\Omega_{1}(E^{\prime}_{1})$ ($\Omega_{2}(E^{\prime}_{2})$) is the number of microstates of first (second) system with energy $E^{\prime}_{1}$ ($E^{\prime}_{2}$), then at this energy distribution, the number of microstates of the composite system would be $\Omega(E^{\prime}_{1})=\Omega_{1}(E^{\prime}_{1})\Omega_{2}(E-E^{\prime}_{1})$. Here we assume that the two systems can independently be in any of their accessible microstates (this assumption is related to our assumption that the interaction energy between the two systems can be ignored). The total number of microstates of the composite system with energy $E$ is $M(E)=\sum_{E^{\prime}_{1}}\Omega(E^{\prime}_{1})$, where the sum is over all possible $E^{\prime}_{1}$ values. As per the premise, the composite system in equilibrium is equally likely to be in any of the $M(E)$ number of microstates. When the systems are thermodynamically large, $M(E)$ will be almost equal to $\Omega(\overline{E}_{1})$, i.e., $M(E)\approx\Omega(\overline{E}_{1})$, where $E^{\prime}_{1}=\overline{E}_{1}$ maximizes the function $\Omega(E^{\prime}_{1})$. This means that the composite system, after settling in a new equilibrium state, will spend most of its time in the microstates corresponding to the energy distribution $\overline{E}_{1}$ and $E-\overline{E}_{1}$ for the first and second systems, respectively.
We stress here that, when the system is in non-equilibrium, it would not be equally probable to be in different microstates (at a given time, a subset of microstates may be more probable than the other subsets). Only in equilibrium, as per the premise, the system is equally probable to be in any of the accessible microstates. Now since in equilibrium the composite system is equally probable to be any of the $M(E)$ microstates, which is approximated by $\Omega(\overline{E}_{1})$, a condition for the new equilibrium between the systems can be established by a maximization condition of $\Omega(E^{\prime}_{1})$ (or, in practice, $\ln\Omega(E^{\prime}_{1})$). After comparing this condition with a standard thermodynamic relation (namely, $\frac{1}{T}=\left(\frac{\partial S}{\partial E}\right)_{N,V}$), one infers that the thermodynamic entropy of the (first) system is equal to the BN entropy, i.e., $k_{B}\ln\Omega_{1}(\overline{E}_{1})$.
From the above discussion, we conclude that the TH entropy and the BN entropy of an isolated system is equivalent ($S_{TH}\equiv S_{BN}$) if the system is large (so that the temperature $T$ can be defined and $M(E)$ can be approximated by $\Omega(\overline{E}_{1})$) and the system is in thermal equilibrium (so that the principle of equal a priori probabilities becomes applicable).
[…]
II.3 Quantum version of statistical entropy
It is not straightforward how one can reinterpret the BN or GB entropy in terms of quantum variable(s). This is because, one of the conditions for equivalence, namely the principle of equal a priori probabilities, needs re-examination. To understand this, we recall that, a classical system moves from one microstate to another due to its time evaluation (each point in a system’s equal-energy trajectory is a microstate and they are statistically equally probable). On the other hand, what is a microstate for a quantum system? It is tempting to think of the eigenstates of the system’s Hamiltonian as the microstates, as one can then refer to all the eigenstates withing a narrow energy window to introduce the principle of equal probabilities for a quantum system. The main problem with this assumption is that the system does not move from one eigenstate to another due to its time evaluation. In the following we discuss how one may resolve this problem.
If we consider an isolated large quantum system with size $N$ (the ‘size’ can be interpreted as the number of particles or constituents of the system), the average gap between any two consecutive eigenvalues will be exponentially small with the system size. We note that the (average) density of states is proportional to the Hilbert space dimension ($D$) and the average gap ($\delta$) is inversely proportional to the density of states. Since $D\propto e^{cN}$, we have $\delta\propto e^{-cN}$, where $c$ is a positive constant. Thus, if $N$ is large enough, it will be practically impossible to prepare the quantum system in a pure state. In other words, because of our said inability, we only can study a system which can be in any of the many closely spaced eigenstates with some probability distribution. Here we may apply the principle of equal a priori probabilities and assume the probability distribution to be uniform in the equilibrium (because, we do not have any specific reason to believe that one state will be preferred over the others in the same energy range).
There are other ways one can justify this equal probabilities principle for a quantum system. For example, due to some (weak) interaction in past, currently the isolated system can be in an entangled state with its environment. This effectively puts the isolated system in a mixed state, i.e., the system can be in any of many microstates (eigenstates) corresponding to a particular energy scale. The associated probability distribution, as per the principle, can be taken to be uniform. Yet another way to justify the principle is the following. Let $H_{0}$ be the main Hamiltonian of the noninteracting system and $H^{\prime}$ represents the effective (weak) interaction among the particles (taken to be perturbation to the main Hamiltonian). Hence the total Hamiltonian describing the system is $H=H_{0}+H^{\prime}$. Now if we take eigenstates of $H_{0}$ to be the microstates, the system will make transitions from one to other microstates during time evaluation due the effect of the interaction $H^{\prime}$. If the effective interaction is taken to be weak, all the eigenstates of $H_{0}$ from a narrow energy window can be assumed to be equally probable, enabling us to use the results of statistical mechanics.
Once we have agreement on quantum version of the principle of equal a priori probabilities, it is not difficult to come up with quantum versions of the statistical entropies. To get the quantum version of the BN (qBN) entropy, we prepare a system in such a way that its energy lies in the narrow energy shell $(E,E+\Delta E]$, where $\Delta E$ is arbitrary but macroscopically small and microscopically large [5]. Let $\{E_{n}\}$ be the set of eigenvalues of the system Hamiltonian and {|$\{\ket{n}\}$ be the corresponding set of eigenkets. As per the premise, the system will be in any state |$\ket{n}$ with equal probability if $E_{n}\in(E,E+\Delta E]$. Subsequently, guided by our understanding of the BN entropy, we can write its quantum version ($S_{qBN}$) in the following way:
$S_{BN}\to S_{qBN}=k_{B}\ln D,$ (1)
where $D$ is the number of eigenvalues within the shell. It may be noted here that one can subdivide the energy shell into different macrostates (subspaces), with one of them having the largest number of microststes (the largest dimension) [7, 5]. For a thermodynamically large system, the dimension of this largest subspace, which corresponds to the thermal equilibrium, is essentially equal to the dimension of the shell ($D$).
To get the quantum version of Gibbs (qGB) entropy, in place of continuous energy, we use energy eigenvalues of the system’s Hamiltonian to calculate the partition function. Here we assume that, the system’s interaction with the reservoir is weak (it had interaction with the reservoir in the past) so that we still can effectively describe the dynamics of the system by its Hamiltonian. As before, let $\{E_{n}\}$ be the set of eigenvalues of the system Hamiltonian. In the spirit of Gibbs distribution for classical system, we can define the probability to find the system with energy $E_{n}$ as $p_{n}=\frac{1}{Q}e^{-\beta E_{n}}$, where $Q$ is the partition function, $Q=\sum_{n}e^{-\beta E_{n}}$. Here $\beta$ denotes the inverse temperature. Subsequently, we can write the quantum version of the GB entropy in the following way:
$S_{GB}\to S_{qGB}=-k_{B}\sum_{n}p_{n}\ln{p_{n}}.$ ) . (2)
III von Neumann Entropy
While reformulating some of the major aspects of statistical mechanics so that it can be used in studying quantum systems, von Neumann provided a formula for entropy of a quantum system in a mixed state [8]. If the density matrix $\rho$ represents the mixed state of a system (not to be confused with the probability distribution function denoted by the same symbol), the entropy of the system, according to von Neumann, is
$S_{VN}=-k_{B}\text{Tr}(\rho\ln\rho).$ (3)
This formula attracted both the applause and criticism alike [10, 11, 12, 13, 14]. In the following two subsections, we first discuss two different perspectives to understand how one can arrive at the formula and what it actually quantifies. Next we highlight some of the major criticisms on the applicability of the formula and explain our views that many of such criticisms can be avoided or are unfounded.
III.1 As basis-independent representation of quantum statistical entropy
One way to understand the VN entropy is to view it as the basis-independent representation of the quantum statistical entropy of a thermodynamic system in equilibrium. The VN entropy for an appropriate density matrix can give us the qBN or qGB entropy. In fact, while deriving the expression of the thermodynamic (TH) entropy of a subsystem, Landau also came up with the same expression appears in Eq. 3 as the basis-independent representation of the quantum statistical entropy [9]. We note that, if we take $\rho=\frac{1}{d}I$, where $d$ is the dimension of the system’s relevant subspace (of the total Hilbert space) and $I$ is the identity matrix, then it is easy to check that $S_{VN}=-k_{B}\text{Tr}(\rho\ln\rho)=-k_{B}\ln d=S_{qBN}$. On the other hand, if we take $\rho=\frac{e^{-\beta H}}{\text{Tr}(e^{-\beta H})}$, where $H$ the system Hamiltonian and $\beta$ is the inverse temperature, then it is difficult to verify that $S_{VN}=S_{qGB}$.
III.2 As entropy of thermodynamic ensemble
Sometimes we may not know in which (pure) state our quantum system is in; von Neumann introduced the concept of the density matrix ($\rho$) to describe the physics of the system [8]. The density matrix is a statistical mixture of many pure states; for example,
$\rho=\sum_{i}p_{i}\ket{\psi_{i}}\bra{\psi_{i}},$ (4)
where the system is expected to be found in state |$\ket{\psi_{i}}$ with probability $p_{i}$ (clearly, $\sum_{i}p_{i}=1$). To calculate the entropy of a system in state $\rho$, von Neumann considered an ensemble, similar to the Gibbsian ensemble for a classical system. The ensemble representing $\rho$ in Eq. 4 will have $\mathcal{N}p_{i}$ number of copies in the state |$\ket{\psi_{i}}$, where $\mathcal{N}$ is the total number of copies in the ensemble.
While following the Gibbsian way, one immediately faces problems in defining entropy of the system in the state $\rho$. First of all, the decomposition in Eq. 4 is not unique and, then, the pure states {|$\{\ket{\psi_{i}}\}$ in a decomposition may not be in general distinguishable (as they are not in general orthonormal). Through thermodynamic considerations, von Neumann proposed a formula (as given in Eq. 3) to quantify entropy of a system in state $\rho$. Without going through the details of his argument, we briefly discuss here the two main features of the VN entropy. First we distinguish two different processes: (1) time evaluation of the system, and (2) measurement on the system. If $\rho(0)$ is the density matrix at time $t=0$ then the density matrix at the $t$ is given by $\rho(t)=U(t)\rho(0)U^{{\dagger}}(t)$, where $U(t)$ is the time evaluation operator. It is not difficult to see that the VN entropy is invariant under this time evaluation process, i.e., $S_{VN}(t)=S_{VN}(0)$. von Neumann showed that a measurement, in general, irreversibly changes the density matrix to another density matrix $\rho\to\rho^{\prime}$, where
$\rho^{\prime}=\sum_{n}\bra{\phi_{n}}\rho\ket{\phi_{n}}P_{n}.$ (5)
We may here note that the set of states {|$\{\ket{\phi_{n}}\}$ are the eigenkets of the observable for which one is performing the measurement. The subtleties of degeneracy can be found in the von Neumann’s work [8]. He showed that the entropy of a system, in general, increases (or remains constant but never decreases) due to the measurement process, i.e., $S_{VN}(\rho^{\prime})\geq S_{VN}(\rho)$. If we identify the first process (time evaluation) and the second process (measurement) with the thermodynamic reversible and irreversible processes, respectively, then these results are consistent with the thermodynamic laws. Before we discuss, in the next subsection, some of the difficulties with the VN entropy formula, we try to intuitively understand the formula below.
As in Sec. II.3, if a pure quantum state is taken to be a microstate, then the qBN entropy of a system in a pure quantum state is zero. In this case, as we see from Eq. 3, the VN entropy is also zero. When the system is not in pure state, rather in a statistical mixture of pure states (represented by $\rho$), a formula for entropy must quantify the amount of “mixedness” (or statistical impurity) in the state $\rho$, similar to how the BN entropy quantifies the amount of “uncertainty” of a system to be found in a microstate. If we now see the decomposition in Eq. 4, it would suggest that the corresponding entropy of the system is $S_{SN}(\{p_{i}\})=-k_{B}\sum_{i}p_{i}\ln p_{i}$, which is the Shannon entropy for the probability distribution $\{p_{i}\}$. Unfortunately, the decomposition of a mixed state is not unique, hence this way of defining entropy is unacceptable. One way to circumvent this problem is to consider all possible decompositions of $\rho$ and take the one for which $S_{SN}(\{p_{i}\})$ is minimum. The advantage of defining entropy this way is that the entropy is now unique for a given $\rho$. In fact it can be proved that the VN entropy of a given $\rho$ corresponds to the aforementioned minimum [15, 16]:
$S_{VN}(\rho)=inf\left\{S_{SN}(\{p_{i}\}):\rho=\sum_{i}p_{i}\ket{\psi_{i}}\bra{\psi_{i}}\right\}.$ | } . (6)
Another way to understand the VN entropy is as follows. If we take a set of orthonormal states to be a set of distinguishable microstates, we can define an entropy from the probabilities associated with these microstates. If {|$\{\ket{\phi_{n}}\}$ is a set of orthonormal states, the entropy one can define is $S_{SN}(\{w_{i}\})=-k_{B}\sum_{n}w_{n}\ln w_{n}$, where $w_{n}=\bra{\phi_{n}}\rho\ket{\phi_{n}}$. Unfortunately, again, this entropy is not unique as it depends on the set of orthonormal states (microstates). To understand how to avoid this problem, we first define a new density matrix $\rho^{\prime}=\sum_{n}w_{n}\ket{\phi_{n}}\bra{\phi_{n}}$ |, where $w_{n}=\bra{\phi_{n}}\rho\ket{\phi_{n}}$. We may note that |$\ket{\phi_{n}}$ is an eigenket of $\rho^{\prime}$ with eigenvalue $w_{n}$. This density matrix ($\rho^{\prime}$) is in the same form as the one defined in Eq. 5. But for this updated density matrix after measurement ($\rho\to\rho^{\prime}$), von Neumann showed that $-\text{Tr}(\rho^{\prime}\ln\rho^{\prime})\geq-\text{Tr}(\rho\ln\rho)$ [8]. Thus we see that the VN entropy corresponds to the minimum of $S_{SN}(\{w_{i}\})$, which we obtain when we choose the microstates as the eigenkets of the given density matrix $\rho$. Accordingly,
$S_{VN}(\rho)=inf\bigg{\{}S_{SN}(\{w_{i}\}):w_{i}=\bra{\phi_{i}}\rho\ket{\phi_{i}}\bigg{\}},$ (7)
where it is understood that $\bra{\phi_{i}}\ket{\phi_{j}}=\delta_{ij}$.
III.3 Criticism against VN entropy: Are we always asking legitimate questions?
Over the century since von Neumann’s work, the entropy after his name found a lot of applications. But along the way it attracted a few major criticisms on its applicability and meaning. Here we address two main objections while trying to interpret the VN entropy as the thermodynamic (TH) entropy.
Time invariance of VN entropy:
The first of these objections is that the VN entropy of an isolated quantum system does not change (more specifically, increase) as the system evolves with time. As seen earlier, $S_{VN}(t)=S_{VN}(0)$ as $\rho(0)\to\rho(t)$ under the time evaluation operator $U(t)$. It is known that, for an isolated system, the TH entropy increases as the system approaches thermal equilibrium. Does not the VN entropy apply to systems in non-equilibrium?
At the outset, we stress that the above issue arises only when the system is not in thermal equilibrium. In the thermal equilibrium, as discussed earlier in Sec. III.1, the VN entropy and the TH entropy are equivalent (under certain general conditions). Before we discuss two different approaches to resolve the problem, we first note that the VN entropy is the quantum-analog of the Gibbs (GB) entropy [17]. Hence, some of the major criticisms faced by the GB entropy are also applicable to the VN entropy. For example, the GB entropy of a system is time invariant since the “shape” of the probability distribution function does not change under Hamiltonian evolution of the system [7]. There are two different ways one can tackle this time-invariance problem. First, one considers that the VN entropy formula needs modification. Accordingly, new entropy formula is proposed which increases as the system evolves with time (this entropy increases and finally saturates in equilibrium). The observational entropy is one such popular alternative to the VN entropy [14, 18]. Second, instead of considering that the VN entropy formula is wrong, one may consider that, while the formula is correct, the time evaluation of density matrix $\rho$ is not correctly described by von Neumann-Liouville equation. We prefer the second option for the following reasons.
Like the TH entropy, the VN entropy is a state function - it depends only on $\rho$ and does not depend on any choice or preference of an observer. The VN entropy formula correctly captures equilibrium thermodynamics of a system. In fact, starting with the VN entropy formula, one can derive the expression of the equilibrium Gibbs distribution using the maximum entropy principle and some other conditions [19]. All these point out that the expression of the VN entropy is a suitable one to study the thermodynamics of a quantum system; main challenge here is to figure out how one the the density matrix evolves from a given initial state to the final one (representing thermal equilibrium). In the present study, we leave out any such detail discussion on how the density matrix evolves.
Subadditivity of VN entropy:
The second major objection against the VN entropy is that it obeys the subadditivity property while the TH entropy is additive (for a large uniform system where the interaction energy between its two parts can be ignored). The subadditivity property is stated as follows: if a system ($AB$) is in state $\rho_{AB}$, and its parts $A$ and $B$ are respectively in states $\rho_{A}$ and $\rho_{B}$ (which are reduced density matrices of the state $\rho_{AB}$), then $S_{VN}(\rho_{AB})\leq S_{VN}(\rho_{A})+S_{VN}(\rho_{B})$ [20]. This implies that, if the system is in a pure state, its VN entropy will be zero but the sum of the VN entropies of its parts is positive (the sum can be zero in a very special case). This contradicts the notion of additivity of the TH entropy. We can circumvent this apparent contradiction in the following way.
As we noted earlier, the VN entropy is the quantum-analog of the GB entropy. The GB entropy is applicable for an open system which is in thermal equilibrium with its bath. For such a thermodynamic system, its TH entropy is equal to the TH entropies of its parts. Now to verify if the VN entropy has this additivity property, we first note that, the quantum thermalization is defined with respect to a smaller subsystem of a large system [21, 22, 1, 5, 23]. Here, a part of the system is thermalized while the rest of the system works as its bath. Therefore, instead of studying additivity property of the whole quantum system, we should study it for a subsystem. It is know that the VN entropy follows a volume law for excited states (especially the states from the middle of its spectrum) [24, 25, 26, 27]. In fact, for our system (see Sec. IV.3), we see that the average VN entropy $\overline{S}_{VN}$ (averaged over all the states from a narrow energy window in the middle of the spectrum) is extensive and grows linearly with subsystem size, as seen in Fig. 1.
Figure 1: In the middle of the spectrum, the average VN entropy ($\overline{S}_{VN}$) grows linearly with subsystem size. Here total system size is $N=16$.
We, therefore, infer here that the VN entropy also obeys the additivity property if interpreted appropriately.
We finish this subsection by concluding that the major criticisms against the VN entropy either can be avoided or can be addressed convincingly if issues are analyzed appropriately. In this regard, we concur with A. Peres when he says that there should be no doubt that von Neumann’s entropy is equivalent to the entropy of classical thermodynamics [28].
IV Entropy of large quantum systems
We already saw in Sec. II and Sec. III.1 that for a thermalized large isolated or open quantum system, the VN entropy is equivalent to the TH entropy (or to an appropriate quantum statistical entropy). It is non-trivial to check this hypothesis for a small subsystem of a large quantum system. In this section we check, mainly through numerical analysis, whether and how the VN entropy is equivalent to the TH entropy for a subsystem. For completeness and more clarity, we first briefly summarize how the VN entropy is equal (or equivalent) to the Th entropy for a thermalized isolated and open system. We exclude from our discussion the systems which are not in equilibrium or the systems which do not thermalize.
Figure 2: The subsystem VN entropy ($S_{VN}$) for individual eigenkets and $\ln(DOS)$ are plotted against the energy of the full system. Here total system size is 16 and subsystem size is 6.
IV.1 Isolated quantum system
As discussed in Sec. II.3, an isolated quantum system will be in a mixed state, for one or multiple reasons. Using the principle of equal a priori probabilities, one can show that the VN entropy is equal to the qBN entropy (which in turn is equivalent to the TH entropy).
Assume that we prepare a system in such a way that its energy lies in the narrow energy shell $(E,E+\Delta E]$, where $\Delta E$ is arbitrary but macroscopically small and microscopically large [5]. Let {|$\{\ket{n}\}$ be the set of the eigenkets of the system Hamiltonian, and $E_{n}$ be the eigenvalue corresponding to |$\ket{n}$. The system will be in any state |$\ket{n}$ with equal probability if $E_{n}\in(E,E+\Delta E]$. Accordingly, the thermodynamic entropy ($S_{TH}$), or equivalently, the quantum Boltmann entropy ($S_{qBN}$) is,
$S_{TH}\equiv S_{qBN}=k_{B}\ln D,$ (8)
where $D$ is the number of eigenvalues within the shell. It may be noted here that one can subdivide the energy shell into different macrostates (subspaces), with one of them having the largest number of microststes (the largest dimension) [7, 5]. For a thermodynamically large system, the dimension of this largest subspace, which corresponds to the thermal equilibrium, is essentially equal to the dimension of the shell ($D$).
Now to calculate the corresponding von Neumann entropy ($S_{VN}$) of the isolated quantum system, we note that the system is in a mixed state, represented by the following density matrix,
$\rho=\sum_{\begin{subarray}{c}n\\ E_{n}\in\left(E,E+\Delta E\right]\end{subarray}}p_{n}\ket{n}\bra{n},$ (9)
where $p_{n}$ is the probability to find the system in state |$\ket{n}$. Since $\Delta E$ is macroscopically small, all the accessible microstates (here, eigenkets) will essentially have almost the same energy. Application of the principle of equal a priori probabilities suggests that $p_{n}=\frac{1}{D}$. As $S_{VN}=-k_{B}\text{Tr}(\rho\ln\rho)$, we need the eigenvalues of $\rho$ to efficiently compute $S_{VN}$. It is clear from Eq. 9 that |$\ket{n}$ is an eigenket of $\rho$ with eigenvalue $\frac{1}{D}$, i.e., $\rho\ket{n}=\frac{1}{D}\ket{n}$ for any energy eigenket from the shell. This implies that
$S_{VN}=-k_{B}\text{Tr}(\rho\ln\rho)=k_{B}\ln D.$ (10)
We see from Eqs. 8 and 10 that, for an isolated quantum system, the VN entropy is the same as the TH entropy ($S_{VN}=S_{qBN}\equiv S_{TH}$), provided that the system is large and in thermal equilibrium. The detail discussion on the conditions under which this equivalence works can be found in Sec. II.2.
IV.2 Weakly interacting open quantum system
Here we consider an open system which is in contact with the thermal bath. Although the system exchanges heat with the bath, we assume that the interaction with the bath is so weak that the energy spectrum of the system is more or less unaffected by the interaction, and the dynamics of the quantum system can still be described (at-least effectively) by its Hamiltonian.
As before, let $\{E_{n}\}$ be the set of eigenvalues of the system Hamiltonian with |$\ket{n}$ being the eigenket corresponding to eigenvalue $E_{n}$. In the spirit of Gibbs distribution for classical system, we can define the probability to find the system with energy $E_{n}$ as $p_{n}=\frac{1}{Q}e^{-\beta E_{n}}$, where $Q$ is the partition function, $Q=\sum_{n}e^{-\beta E_{n}}$. Here $\beta$ denotes the inverse temperature. We can now get the TH entropy, quantified by the quantum Gibbs entropy, as,
$S_{TH}\equiv S_{qGB}=-k_{B}\sum_{n}p_{n}\ln{p_{n}}.$ ) . (11)
Figure 3: The average VN entropy ($\overline{S}_{VN}$) for subsystem is plotted against $\ln(DOS)$ across the energy spectrum (separately for the left and the right halves of the spectrum). Our system size is 16 and subsystem size is 6.
To find the VN entropy of the open system, one takes the following density operator: $\rho=\frac{e^{-\beta H}}{Tr(e^{-\beta H})}$, where the inverse temperature can be fixed by solving the equation $Tr(He^{-\beta H})=U$ with $U$ being the (given) internal or average energy. In fact, for overwhelmingly large number of the pure states (of the system-bath combination), it can be quite generally shown that, a weakly interacting open quantum system is very well described by a (reduced) density matrix which is equivalent to the known Gibbsian distribution [29, 30]. As a result, the VN entropy of the open system is,
$S_{VN}=-k_{B}\text{Tr}(\rho\ln\rho)=-k_{B}\sum_{n}p_{n}\ln{p_{n}},$ (12)
where the expression of $p_{n}$ is the same as the one appears in the formula for $S_{qGB}$ in Eq. 11. We now see from Eqs. 11 and 12 that the the VN entropy is the same as the TH entropy ($S_{VN}=S_{qBN}\equiv S_{TH}$) for a weakly interacting open quantum system in thermal equilibrium. The detail discussion on the conditions under which this equivalence works can be found in Sec. II.2.
IV.3 Subsystem of a quantum system
As mentioned in the beginning of this section (Sec. IV), it is non-trivial to verify if the VN entropy is equivalent to the TH entropy for a subsystem (of a much larger quantum system in thermal equilibrium). We analyze this possibility here.
We consider a large composite quantum system $AB$ in a state described by the density matrix $\rho$. We also assume that the subsystem $A$ is macroscopically large but much smaller compare to its counterpart $B$. We here show, mainly by the numerical analysis, that the VN entropy of the subsystem $A$ is proportional to its TH entropy.
To first calculate the TH entropy of a subsystem, we again assume that the total system is prepared with an energy lying in $(E,E+\Delta E]$. As discussed earlier, the system will be in a mixed state $\rho$, as appear in Eq. 9. The TH entropy (qBN entropy) of the full system is $S_{TH}\equiv S_{qBN}=k_{B}\ln D$, as in Eq. 8, where the $D$ is the number of eigenvalues of the system lying in the aforementioned energy shell. Now since the TH entropy is additive, we can assume that the TH entropy of the subsystem (denoted as $S^{sb}_{TH}$ ) is proportion to $k_{B}\ln D$ or to the $k_{B}\ln(DOS)$, where $DOS$ is the density-of-states of the full system at the energy scale $E$. Accordingly,
$S^{sb}_{TH}\propto k_{B}\ln(DOS),$ (13)
where the proportionality constant depends on the size of the subsystem in comparison to the size of the full system.
Now to calculate the VN entropy of the subsystem, we first need to calculate the reduced density matrix of the subsystem $\rho_{A}=\text{Tr}_{B}(\rho)$, where $\rho$ is the density matrix of the full system, as given in Eq. 9 (with $p_{n}=\frac{1}{D}$). Denoting the reduced density matrix corresponding to the eigenket |$\ket{n}$ by $\rho_{A}^{(n)}$, where $\rho_{A}^{(n)}=\text{Tr}_{B}(\ket{n}\bra{n})$, we have
$\rho_{A}=\frac{1}{D}\sum_{n}\rho_{A}^{(n)}.$ (14)
Since the subsystem is in the mixed state $\rho_{A}$, we now can calculate the VN entropy of the subsystem directly from $\rho_{A}$, i.e,
$S^{sb}_{VN}=-k_{B}\text{Tr}(\rho_{A}\ln\rho_{A}),$ (15)
where $\rho_{A}$ is given by Eq. 14. For a system which thermalizes (or obeys the eigenstate thermalization hypothesis [21, 22]), we can also have
$S^{sb}_{VN}=-k_{B}\text{Tr}(\rho_{A}\ln\rho_{A})\approx-\frac{k_{B}}{D}\sum_{n}\text{Tr}(\rho_{A}^{(n)}\ln\rho_{A}^{(n)}).$ (16)
This approximate equality works since, as per the hypothesis, the expectation value of a local operator (say, $O$) is almost the same for different eigenkets from a narrow energy window, i.e., $\text{Tr}(\rho_{A}^{(n)}O)\approx\text{Tr}(\rho_{A}^{(m)}O)$. This is possible if the set of eigenvalues of the reduced density matrix $\rho_{A}^{(n)}$ is almost the same for different $n$. This implies that $\text{Tr}(\rho_{A}^{(n)}\ln\rho_{A}^{(n)})\approx\text{Tr}(\rho_{A}^{(m)}\ln\rho_{A}^{(m)})$, which guarantees the approximate equality in Eq. 16. If we write $\overline{S}_{VN}=-\frac{k_{B}}{D}\sum_{n}\text{Tr}(\rho_{A}^{(n)}\ln\rho_{A}^{(n)})$, then $S^{sb}_{VN}\approx\overline{S}_{VN}$ for a thermalized system.
We note that, if the system does not thermalize (or is integrable), then the VN entropy for the individual eigenkets ($-k_{B}\text{Tr}(\rho_{A}^{(n)}\ln\rho_{A}^{(n)})$) can take widely different values. This in fact can be seen in Fig. 2, where the subsystem VN entropy for different eigenkets are plotted against the energy across the spectrum. We here mention that, for all our numerical results, we take the following one-dimensional (with open boundary condition) Heisenberg spin-1/2 chain with the next-nearest neighbor interactions:
$H=J\left(\sum_{i=1}^{N-1}\vec{S}_{i}\cdot\vec{S}_{i+1}+\Delta_{2}\sum_{i=1}^{N-2}S^{z}_{i}S^{z}_{i+2}\right).$ (17)
In our calculations, we take $J=1$ and the size of the system $N=16$. This model system is integrable when $\Delta_{2}=0$ and is non-integrable (hence thermalizes) when $\Delta_{2}>0$ [31].
To make more sense of the results in Fig. 2, we plot the subsystem’s average VN entropy ($\overline{S}_{VN}$) against $\ln(DOS)$ (separately for the left and right halves of the spectrum). The results are shown in Fig. 3. We find that, for a thermalized system ($\Delta_{2}>0$), its subsystem’s average VN entropy ($\overline{S}_{VN}$) is proportional to $\ln(DOS)$ with a proportionality constant approximately equal to $\frac{N_{sb}}{N}$ (in our calculations, the subsystem size $N_{sb}=6$ and the full system size $N=16$). This proportionality constant takes a different value (a smaller one) if the system is not thermalized. We take $k_{B}=1$ for all our numerical calculations of entropy. Thus our numerical results suggest that,
$S^{sb}_{VN}\approx\overline{S}_{VN}\propto\ln(DOS).$ (18)
From Eqs. 13 and 18, we conclude that, like for isolated and open systems, for a subsystem also the VN entropy is equivalent to the TH entropy (qBN entropy), i.e., $S^{sb}_{VN}\equiv S^{sb}_{TH}$.
It may be worth mentioning here that, the result that the subsystem’s average VN entropy is proportional to system’s $\ln(DOS)$, is earlier reported in [32] in a different context. We also note that, elsewhere [33], the thermodynamic entropy of a subsystem is defined as what we call here $S^{sb}_{VN}$. As discussed in this work, thermodynamic (TH) entropy is appropriately calculated as the logarithm of $DOS$ (since, $S_{TH}\equiv S_{qBN}$ under some general conditions).
V Conclusion
In this work we analyze whether and when the von Neumann (VN) entropy is equal or equivalent to the thermodynamic (TH) entropy. In reviewing the major objections about the VN entropy, we explain that the concerns about the time-invariance and subadditivity properties can be addressed convincingly. We show in this study that, for large thermalized isolated or open system, the VN entropy is equivalent to the TH entropy. Next, through numerical analysis, we also show that the VN entropy of a subsystem of a large thermalized system is equivalent to its TH entropy (qBN entropy).

### v2

@@
 The debate whether the von Neumann (VN) entropy is a suitable quantum version of the thermodynamic (TH) entropy has a long history.
-In this regard, we briefly review some of the main reservations about the VN entropy and explain that the objections about its time-invariance and subadditivity properties can be either avoided or addressed convincingly.
-In a broader context, we analyze here whether and when the VN entropy is the same or equivalent to the TH entropy.
-For a thermalized isolated or open system, the VN entropy for an appropriately chosen density operator is the same as the quantum version of, respectively, the Boltzmann or Gibbs entropy (these latter entropies are equivalent to the TH entropy for large thermalized systems).
-Since the quantum thermalization is essentially defined for a subsystem of a much larger system, it is important to investigate if the VN entropy of a subsystem is equivalent to its TH entropy.
-We here show that the VN entropy of a small subsystem of a large thermalized system is proportional to the quantum statistical entropy (Boltzmann’s version) of the subsystem.
-For relevant numerical results, we take a one-dimensional spin-1/2 chain with next-nearest neighbor interactions.
+In this regard, we briefly review some of the main reservations about the VN entropy and explain that the objections about its time-invariance and subadditivity properties can be addressed convincingly.
+In a broader context, we here analyze in detail whether and when the VN entropy is the same or equivalent to the TH entropy.
+We find that the VN entropy of a large isolated system is equivalent to its TH entropy if the system is in a mixed state described by a microcanonical (MC) ensemble.
+This equivalence also holds for open system and subsystem if the full system (i.e.
+system-bath or subsystem-bath combination) is in a mixed state described by an MC ensemble or in a typical pure state (or if the Eigenstate Thermalization Hypothesis, ETH, holds for the full system).
+Unlike isolated and open system, demonstrating the equivalence for a subsystem is not straightforward.
+We take a one-dimensional spin-1/2 model to numerically show that, if the full system is in a pure state, then the VN entropy of a subsystem is equivalent to its TH entropy if the ETH holds for the full system.
+We also demonstrate that this equivalence always works (irrespective of whether the ETH holds) if the full system is in a mixed state describe by an MC ensemble.
 preprint: APS/123-QED
@@
 Boltzmann, and later Gibbs, showed us how the TH entropy can be calculated from the microscopic details using statistical approaches [4, 6].
-For isolated system, Boltzmann gave formula (actually Planck first wrote down the explicit expression for the Boltzmann entropy) for the TH entropy in terms of the volume of the relevant phase space (or the number of associated microstates).
-For the open system, Gibbs on the other hand, using his ensemble theory, provided the formula for the TH entropy in terms for probability distribution function in phase space.
-The quantum versions of these statistical entropies (Boltzmann’s and Gibbs’) are somewhat intuitively clear [7, 5]: for an isolated quantum system, the entropy is given in terms of the dimension of the relevant subspace of the system’s Hilbert space while for an open quantum system, the entropy can be calculated from the partition function (which depends on the eigenvalues of the appropriate Hamiltonian of the system).
-Question, however, remains on the quantum version of the TH entropy for a part (subsystem) of a quantum system.
-We will address this issue separately in this work.
-From the thermodynamical considerations, von Neumann (VN) gave a formula for entropy of a quantum ensemble (representing a given mixed state) [8].
-This entropy, known as VN entropy, attracted a lot of attention and criticism in regards to its connection to the TH entropy.
-We show in this work that, for isolated and open quantum system or for small part of a large quantum system, the VN entropy is equivalent to the TH entropy if the quantum system is large and in thermal equilibrium.
-We largely exclude from our discussions the quantum systems which do not thermalize or systems not in equilibrium.
-It is worth mentioning here the logical steps we follow in this work while establishing the said equivalence between the entropies and finding the conditions under which the equivalence works.
-First, identify under what conditions the TH entropy is equivalent to the statistical entropy (either Boltzmann’s or Gibbs’ version), then reinterpret the statistical entropy in terms of the quantum variables (in Boltzmann’s case: dimension of relevant subspace in place of volume of accessible phase-space, and in Gibbs’ case: energy eigenvalues in place of continuous classical energy).
-In the last step of understanding the equivalence, one checks if the VN entropy of an appropriate density matrix is equivalent to the quantum version of the statistical entropy (i.e.
+For an isolated system, Boltzmann gave a formula (actually Planck first wrote down the explicit expression for the Boltzmann entropy) for the TH entropy in terms of the volume of the relevant phase space (or the number of associated microstates).
+For the open system, Gibbs on the other hand, using his ensemble theory, provided the formula for the TH entropy in terms of an appropriate probability distribution function in phase space.
+Although the quantum versions of these statistical entropies for the isolated and open systems in equilibrium are somewhat intuitively clear [7, 5], one needs convincing justifications, which we discuss in Sec.
+II.
+The question, however, remains on the quantum version of the TH entropy for a part (subsystem) of a quantum system.
+In this work, we will discuss this issue in detail.
+As noted, for a quantum system in thermal equilibrium, writing down the quantum version of the Gibbs (GB) entropy is somewhat easy (see, Secs.
+II.3 and III.1).
+But when a quantum system is in a generic mixed state and not in equilibrium, what would be the value of entropy?
+From thermodynamical considerations, von Neumann (VN) gave a formula for the entropy of a quantum ensemble (representing a given mixed state) [8].
+This entropy, known as VN entropy, attracted a great deal of attention and a lot of criticism in regards to its connection to the TH entropy.
+We show in this work that, for isolated and open quantum system or for a subsystem of a (much) larger quantum system, the VN entropy is equivalent to the TH entropy if the quantum system (or subsystem) is large and in thermal equilibrium.
+We largely exclude from our discussions the quantum systems that do not thermalize or systems that are not in equilibrium.
+We here, though, deliberate on two main criticisms on the VN entropy, related to the time-invariance and subadditivity, and argue that these criticisms can be avoided or addressed convincingly.
+We outline here the logical steps we follow in this work while establishing the said equivalence between the VN and TH entropies and finding the conditions under which the equivalence works.
+First, we identify under what conditions the TH entropy is equivalent to the statistical entropy (either Boltzmann’s or Gibbs’ version), then reinterpret the statistical entropy in terms of the quantum variables (in Boltzmann’s case: the dimension of relevant subspace in place of the volume of accessible phase-space, and in Gibbs’ case: the energy eigenvalues in place of the continuous classical energy).
+In the last step of establishing the equivalence, we check if the VN entropy of an appropriate density matrix is equivalent to the quantum version of the statistical entropy (i.e.
 the reinterpreted statistical entropy in terms of quantum variables).
-To better understand the equivalence, one should also check if the the VN entropy inherits the important properties of the equilibrium TH entropy.
-Accordingly, whereever appropriate, we briefly deliberate, if the VN entropy is a state function, additive and whether it increases for isolated system due to an irreversible process.
+To better understand the equivalence, one should also check if the VN entropy inherits the important properties of the equilibrium TH entropy.
+Accordingly, wherever appropriate, we briefly deliberate if the VN entropy is a state function, has the additivity property and whether it increases for isolated system due to an irreversible process.
 We organize our paper in the following way.
 In the next section (Sec.
-II), we review the thermodynamic and statistical entropies, and subsequently, discuss the suitable quantum versions of these statistical entropies.
+II), we review the thermodynamic and statistical entropies, and discuss the condition of equivalence between these entropies.
+This is followed by a discussion on the suitable quantum versions of the statistical entropies.
 In Sec.
-III, we discuss von Neumann entropy and some reservations about the entropy while it comes to representing the thermodynamic entropy.
+III, we discuss von Neumann entropy and some reservations about the entropy when it comes to representing the thermodynamic entropy.
 Next, in Sec.
@@
 Here the equality sign works only for a reversible cycle and this helps us to define a thermodynamic (TH) state function, called entropy.
-Streaky speaking, this state function (thermodynamic entropy, $S_{TH}$) is well defined only for a thermodynamic system in equilibrium.
+Strictly speaking, this state function (thermodynamic entropy, $S_{TH}$) is well defined only for a thermodynamic system in equilibrium.
 If a system in equilibrium at temperature $T$ receives $d\bar{}\hskip 1.00006ptq$ amount of heat reversibly and subsequently reaches an equilibrium again, then the difference in entropy between these two equilibrium phases would be $dS_{TH}=\frac{d\bar{}\hskip 0.70004ptq}{T}$.
+In contrast, if the process is irreversible, then $dS_{TH}>\frac{d\bar{}\hskip 0.70004ptq}{T}$.
+This shows, for any isolated system undergoing a process ($d\bar{}\hskip 1.00006ptq=0$), the change in its entropy is $dS_{TH}\geq 0$.
 The Boltzmann’s version of statistical entropy is defined by the number of accessible microstates $\Omega$ of a system (with a fixed energy).
@@
 In the following subsection we discuss under what conditions the TH entropy and the statistical entropy (BN or GB entropy) are equivalent.
-II.2 Conditions of equivalence between thermodynamic and statistical entropy
+II.2 Conditions of equivalence between thermodynamic and statistical entropies
 To understand the conditions under which the equivalence works, we outline here a method to establish such equivalence.
@@
 Before the contact, the systems were in thermal equilibrium and after the contact they will again come to a common equilibrium via exchange of heat.
-Although the systems exchange heat, total energy ($E=E_{1}+E_{2}$) of the composite system remains constant; here one ignores the interaction energy between the systems.
-In the physical setup to establish the equivalence, the main premise is that a thermodynamic system in equilibrium with a given energy can be in any of a certain number of microstates with an equal probability (the principle of equal a priori probabilities).
-Here the number of accessible microstates is a function of the system’s energy.
-If $\Omega_{1}(E^{\prime}_{1})$ ($\Omega_{2}(E^{\prime}_{2})$) is the number of microstates of first (second) system with energy $E^{\prime}_{1}$ ($E^{\prime}_{2}$), then at this energy distribution, the number of microstates of the composite system would be $\Omega(E^{\prime}_{1})=\Omega_{1}(E^{\prime}_{1})\Omega_{2}(E-E^{\prime}_{1})$.
-Here we assume that the two systems can independently be in any of their accessible microstates (this assumption is related to our assumption that the interaction energy between the two systems can be ignored).
-The total number of microstates of the composite system with energy $E$ is $M(E)=\sum_{E^{\prime}_{1}}\Omega(E^{\prime}_{1})$, where the sum is over all possible $E^{\prime}_{1}$ values.
+Although the systems exchange heat, the total energy ($E=E_{1}+E_{2}$) of the composite system remains constant; here one ignores the interaction energy between the systems.
+In the physical setup to establish the equivalence, one generally employs the principle of equal a priori probabilities.
+According to this premise, a thermodynamic system in equilibrium with a given energy can be found in any of the available microstates with equal probability.
+Here the number of available microstates is a function of the system’s energy.
+If $\Omega_{1}(E^{\prime}_{1})$ ($\Omega_{2}(E^{\prime}_{2})$) is the number of microstates of first (second) system with energy $E^{\prime}_{1}$ ($E^{\prime}_{2}$), then in this energy distribution, the number of microstates of the composite system would be $\Omega(E^{\prime}_{1})=\Omega_{1}(E^{\prime}_{1})\Omega_{2}(E-E^{\prime}_{1})$.
+We here assume that the two systems can independently be in any of their available microstates (this assumption is related to our assumption that the interaction energy between the two systems can be ignored).
+The total number of microstates of the composite system with energy $E$ is $M(E)=\sum_{E^{\prime}_{1}}\Omega(E^{\prime}_{1})$, where the sum is over all possible $E^{\prime}_{1}$ values (i.e., over all possible energy distributions).
 As per the premise, the composite system in equilibrium is equally likely to be in any of the $M(E)$ number of microstates.
@@
 Only in equilibrium, as per the premise, the system is equally probable to be in any of the accessible microstates.
-Now since in equilibrium the composite system is equally probable to be any of the $M(E)$ microstates, which is approximated by $\Omega(\overline{E}_{1})$, a condition for the new equilibrium between the systems can be established by a maximization condition of $\Omega(E^{\prime}_{1})$ (or, in practice, $\ln\Omega(E^{\prime}_{1})$).
-After comparing this condition with a standard thermodynamic relation (namely, $\frac{1}{T}=\left(\frac{\partial S}{\partial E}\right)_{N,V}$), one infers that the thermodynamic entropy of the (first) system is equal to the BN entropy, i.e., $k_{B}\ln\Omega_{1}(\overline{E}_{1})$.
-From the above discussion, we conclude that the TH entropy and the BN entropy of an isolated system is equivalent ($S_{TH}\equiv S_{BN}$) if the system is large (so that the temperature $T$ can be defined and $M(E)$ can be approximated by $\Omega(\overline{E}_{1})$) and the system is in thermal equilibrium (so that the principle of equal a priori probabilities becomes applicable).
+Now since in equilibrium, the composite system is equally probable to be found in any of the $M(E)$ microstates, which is approximated by $\Omega(\overline{E}_{1})$, a condition for the new equilibrium between the systems can be established by a maximization condition of $\Omega(E^{\prime}_{1})$ (or, in practice, $\ln\Omega(E^{\prime}_{1})$).
+After comparing this condition with a standard thermodynamic relation (namely, $\frac{1}{T}=\left(\frac{\partial S_{TH}}{\partial E}\right)_{N,V}$), one infers that the thermodynamic entropy of the (first) system is equal to the BN entropy, i.e., $k_{B}\ln\Omega_{1}(\overline{E}_{1})$.
+From the above discussion, we conclude that the TH entropy and the BN entropy of an isolated system are equivalent ($S_{TH}\equiv S_{BN}$) if the system is large (so that the temperature $T$ can be defined and $M(E)$ can be approximated by $\Omega(\overline{E}_{1})$) and the system is in thermal equilibrium (so that the principle of equal a priori probabilities becomes applicable).
 […]
 II.3 Quantum version of statistical entropy
-It is not straightforward how one can reinterpret the BN or GB entropy in terms of quantum variable(s).
-This is because, one of the conditions for equivalence, namely the principle of equal a priori probabilities, needs re-examination.
+Intuitively it is not very difficult to guess the quantum version of a statistical entropy.
+However, it is not straightforward to see how the conditions discussed in the preceding subsection are satisfied for a quantum system.
+For example, one of the conditions for equivalence, namely the principle of equal a priori probabilities, needs re-examination.
 To understand this, we recall that, a classical system moves from one microstate to another due to its time evaluation (each point in a system’s equal-energy trajectory is a microstate and they are statistically equally probable).
@@
 In the following we discuss how one may resolve this problem.
-If we consider an isolated large quantum system with size $N$ (the ‘size’ can be interpreted as the number of particles or constituents of the system), the average gap between any two consecutive eigenvalues will be exponentially small with the system size.
+If we consider an isolated large quantum system with size $N$ (the ‘size’ can be interpreted as the number of particles in or constituents of the system), the average gap between any two consecutive eigenvalues will be exponentially small with the system size.
 We note that the (average) density of states is proportional to the Hilbert space dimension ($D$) and the average gap ($\delta$) is inversely proportional to the density of states.
@@
 Hence the total Hamiltonian describing the system is $H=H_{0}+H^{\prime}$.
-Now if we take eigenstates of $H_{0}$ to be the microstates, the system will make transitions from one to other microstates during time evaluation due the effect of the interaction $H^{\prime}$.
-If the effective interaction is taken to be weak, all the eigenstates of $H_{0}$ from a narrow energy window can be assumed to be equally probable, enabling us to use the results of statistical mechanics.
-Once we have agreement on quantum version of the principle of equal a priori probabilities, it is not difficult to come up with quantum versions of the statistical entropies.
+Now if we take the eigenstates of $H_{0}$ to be the microstates, then the system will make transitions from one to other microstates during time evaluation due to the effect of the interaction $H^{\prime}$.
+If the effective interaction is taken to be weak, all the eigenstates of $H_{0}$ from a narrow energy window can be assumed to be equally probable, allowing us to use the results of statistical mechanics.
+Once we have agreement on the quantum version of the principle of equal a priori probabilities, it is not difficult to come up with quantum versions of the statistical entropies.
 To get the quantum version of the BN (qBN) entropy, we prepare a system in such a way that its energy lies in the narrow energy shell $(E,E+\Delta E]$, where $\Delta E$ is arbitrary but macroscopically small and microscopically large [5].
@@
 For a thermodynamically large system, the dimension of this largest subspace, which corresponds to the thermal equilibrium, is essentially equal to the dimension of the shell ($D$).
-To get the quantum version of Gibbs (qGB) entropy, in place of continuous energy, we use energy eigenvalues of the system’s Hamiltonian to calculate the partition function.
-Here we assume that, the system’s interaction with the reservoir is weak (it had interaction with the reservoir in the past) so that we still can effectively describe the dynamics of the system by its Hamiltonian.
-As before, let $\{E_{n}\}$ be the set of eigenvalues of the system Hamiltonian.
-In the spirit of Gibbs distribution for classical system, we can define the probability to find the system with energy $E_{n}$ as $p_{n}=\frac{1}{Q}e^{-\beta E_{n}}$, where $Q$ is the partition function, $Q=\sum_{n}e^{-\beta E_{n}}$.
+To get the quantum version of the Gibbs (qGB) entropy, in place of continuous classical energy, we use energy eigenvalues of the system’s Hamiltonian to calculate the partition function.
+Here we assume that the system’s interaction with the reservoir is weak (or it had interaction with the reservoir in the past), so that we still can effectively describe the dynamics of the system by its Hamiltonian.
+As before, let $\{E_{n}\}$ be the set of eigenvalues of the Hamiltonian of the (open) system.
+In the spirit of the Gibbs distribution for the classical system, we can define the probability to find the system with energy $E_{n}$ as $p_{n}=\frac{1}{Q}e^{-\beta E_{n}}$, where $Q$ is the partition function, $Q=\sum_{n}e^{-\beta E_{n}}$.
 Here $\beta$ denotes the inverse temperature.
@@
 $S_{GB}\to S_{qGB}=-k_{B}\sum_{n}p_{n}\ln{p_{n}}.$ ) . (2)
+We here briefly note that the probabilities $p_{n}$ are the eigenvalues of the density matrix $\rho_{\beta}=\frac{e^{-\beta H}}{\text{Tr}(e^{-\beta H})}$, where $H$ is the Hamiltonian of the (open) system.
+Using the density matrix $\rho_{\beta}$, the quantum Gibbs (qGB) entropy in Eq.2 can also be expressed in a basis-independent form, as discussed in Sec.III.1.
+It may be emphasized that the relevant density matrix is found here indirectly: starting from a classical (open) system, we first find the probability distribution and then reinterpreting the quantities in terms of the quantum variables, we get the density matrix.
+Instead, it is also possible to get the density matrix $\rho_{\beta}$ directly, starting with a quantum system [10, 11].
+In this work, we prefer the first method since it directly reveals when and how the TH entropy is equivalent to the qGB entropy.
 III von Neumann Entropy
@@
 This formula attracted both the applause and criticism alike [12, 13, 14, 15, 16].
-In the following two subsections, we first discuss two different perspectives to understand how one can arrive at the formula and what it actually quantifies.
-Next we highlight some of the major criticisms on the applicability of the formula and explain our views that many of such criticisms can be avoided or are unfounded.
+In the following subsections, we first discuss two different perspectives to understand how one can arrive at the formula and what it actually quantifies.
+Next we highlight some of the major criticisms on the applicability of the formula and explain our views that many of such criticisms can be avoided or can be addressed convincingly.
 III.1 As basis-independent representation of quantum statistical entropy
@@
 The VN entropy for an appropriate density matrix can give us the qBN or qGB entropy.
-In fact, while deriving the expression of the thermodynamic (TH) entropy of a subsystem, Landau also came up with the same expression appears in Eq.
+In fact, while deriving the expression of the thermodynamic (TH) entropy of a subsystem, Landau also came up with the same expression that appears in Eq.
 3 as the basis-independent representation of the quantum statistical entropy [9].
-We note that, if we take $\rho=\frac{1}{d}I$, where $d$ is the dimension of the system’s relevant subspace (of the total Hilbert space) and $I$ is the identity matrix, then it is easy to check that $S_{VN}=-k_{B}\text{Tr}(\rho\ln\rho)=-k_{B}\ln d=S_{qBN}$.
-On the other hand, if we take $\rho=\frac{e^{-\beta H}}{\text{Tr}(e^{-\beta H})}$, where $H$ the system Hamiltonian and $\beta$ is the inverse temperature, then it is difficult to verify that $S_{VN}=S_{qGB}$.
+For an isolated system, described by a microcanonical (mc) ensemble, we get the following from Eq.
+1: $S_{qBN}=-k_{B}\ln D=-k_{B}\text{Tr}(\rho_{mc}\ln\rho_{mc})=S_{VN}$, where $\rho_{mc}=\frac{1}{D}I$ with $D$ being the dimension of the relevant subspace of the Hilbert space of the system and $I$ is the identity matrix.
+So we find here that the VN entropy can be viewed as a basis-independent representation of the qBN entropy of an isolated system in thermal equilibrium.
+Similarly, for an open system, we get the following from the Eq.
+2, $S_{qGB}=-k_{B}\sum_{n}p_{n}\ln{p_{n}}=-k_{B}\text{Tr}(\rho_{\beta}\ln\rho_{\beta})=S_{VN}$, where $\rho_{\beta}=\frac{e^{-\beta H}}{\text{Tr}(e^{-\beta H})}$.
+Here $H$ and $\beta$ are, respectively, the Hamiltonian and the inverse temperature of the open system.
+We note again that the VN entropy of an open system in equilibrium can be viewed as a basis-independent representation of the qGB entropy.
 III.2 As entropy of thermodynamic ensemble
-Sometimes we may not know in which (pure) state our quantum system is in; von Neumann introduced the concept of the density matrix ($\rho$) to describe the physics of the system [8].
+Sometimes we may not know in which pure state our quantum system is in; von Neumann introduced the concept of the density matrix ($\rho$) to describe the physics of such system [8].
 The density matrix is a statistical mixture of many pure states; for example,
@@
 Through thermodynamic considerations, von Neumann proposed a formula (as given in Eq.
-3) to quantify entropy of a system in state $\rho$.
-Without going through the details of his argument, we briefly discuss here the two main features of the VN entropy.
-First we distinguish two different processes: (1) time evaluation of the system, and (2) measurement on the system.
+3) to quantify the entropy of a system in state $\rho$.
+We will not go through his arguments here; rather, we will discuss two alternative ways to understand the VN entropy.
+Before that, we briefly discuss below the two main quantum processes and their effects on the VN entropy.
+As discussed by von Neumann, we can consider two basic processes that a system may undergo: (1) time evaluation of the system, and (2) measurement on the system.
 If $\rho(0)$ is the density matrix at time $t=0$ then the density matrix at the $t$ is given by $\rho(t)=U(t)\rho(0)U^{{\dagger}}(t)$, where $U(t)$ is the time evaluation operator.
@@
 $\rho^{\prime}=\sum_{n}\bra{\phi_{n}}\rho\ket{\phi_{n}}P_{n}.$ (5)
-We may here note that the set of states {|$\{\ket{\phi_{n}}\}$ are the eigenkets of the observable for which one is performing the measurement.
-The subtleties of degeneracy can be found in the von Neumann’s work [8].
+We note that the set of states {|$\{\ket{\phi_{n}}\}$ are the eigenkets of the observable for which one is performing the measurement and the projection operator $P_{n}=\ket{n}\bra{n}$ |.
+The subtleties in dealing with degeneracy can be found in von Neumann’s work [8].
 He showed that the entropy of a system, in general, increases (or remains constant but never decreases) due to the measurement process, i.e., $S_{VN}(\rho^{\prime})\geq S_{VN}(\rho)$.
 If we identify the first process (time evaluation) and the second process (measurement) with the thermodynamic reversible and irreversible processes, respectively, then these results are consistent with the thermodynamic laws.
-Before we discuss, in the next subsection, some of the difficulties with the VN entropy formula, we try to intuitively understand the formula below.
+In the following, we now discuss two different ways one can intuitively understand the entropy formula.
 As in Sec.
-II.3, if a pure quantum state is taken to be a microstate, then the qBN entropy of a system in a pure quantum state is zero.
+II.3, if a pure quantum state is taken to be a microstate, then the qBN entropy of a system in a pure quantum state is zero (see Eq.
+1).
 In this case, as we see from Eq.
@@
 If we now see the decomposition in Eq.
-4, it would suggest that the corresponding entropy of the system is $S_{SN}(\{p_{i}\})=-k_{B}\sum_{i}p_{i}\ln p_{i}$, which is the Shannon entropy for the probability distribution $\{p_{i}\}$.
+4, it would suggest that the corresponding entropy of the system is $S_{SN}(\{p_{i}\})=-k_{B}\sum_{i}p_{i}\ln p_{i}$, which is the Shannon (SN) entropy for the probability distribution $\{p_{i}\}$.
 Unfortunately, the decomposition of a mixed state is not unique, hence this way of defining entropy is unacceptable.
@@
 5.
-But for this updated density matrix after measurement ($\rho\to\rho^{\prime}$), von Neumann showed that $-\text{Tr}(\rho^{\prime}\ln\rho^{\prime})\geq-\text{Tr}(\rho\ln\rho)$ [8].
+For this updated density matrix after measurement ($\rho\to\rho^{\prime}$), as mentioned earlier, von Neumann showed that $-\text{Tr}(\rho^{\prime}\ln\rho^{\prime})\geq-\text{Tr}(\rho\ln\rho)$ [8].
 Thus we see that the VN entropy corresponds to the minimum of $S_{SN}(\{w_{i}\})$, which we obtain when we choose the microstates as the eigenkets of the given density matrix $\rho$.
@@
 where it is understood that $\bra{\phi_{i}}\ket{\phi_{j}}=\delta_{ij}$.
-III.3 Criticism against VN entropy: Are we always asking legitimate questions?
+III.3 Addressing objections against VN entropy
 Over the century since von Neumann’s work, the entropy after his name found a lot of applications.
@@
 Here we address two main objections while trying to interpret the VN entropy as the thermodynamic (TH) entropy.
-Time invariance of VN entropy:
-The first of these objections is that the VN entropy of an isolated quantum system does not change (more specifically, increase) as the system evolves with time.
-As seen earlier, $S_{VN}(t)=S_{VN}(0)$ as $\rho(0)\to\rho(t)$ under the time evaluation operator $U(t)$.
+Time-invariance of VN entropy:
+The first of these objections is that the VN entropy of an isolated quantum system does not change (more specifically, increase) as the system evolves (unitarily) with time.
+As seen earlier, $S_{VN}(t)=S_{VN}(0)$ when $\rho(0)\to\rho(t)$ under the time evaluation operator $U(t)$.
 It is known that, for an isolated system, the TH entropy increases as the system approaches thermal equilibrium.
 Does not the VN entropy apply to systems in non-equilibrium?
-At the outset, we stress that the above issue arises only when the system is not in thermal equilibrium.
+At the outset, we stress that the above issue arises only when the system is in a non-equilibrium state and we want to know the entropy during a process.
 In the thermal equilibrium, as discussed earlier in Sec.
 III.1, the VN entropy and the TH entropy are equivalent (under certain general conditions).
-Before we discuss two different approaches to resolve the problem, we first note that the VN entropy is the quantum-analog of the Gibbs (GB) entropy [17].
-Hence, some of the major criticisms faced by the GB entropy are also applicable to the VN entropy.
-For example, the GB entropy of a system is time invariant since the “shape” of the probability distribution function does not change under Hamiltonian evolution of the system [7].
-There are two different ways one can tackle this time-invariance problem.
-First, one considers that the VN entropy formula needs modification.
-Accordingly, new entropy formula is proposed which increases as the system evolves with time (this entropy increases and finally saturates in equilibrium).
-The observational entropy is one such popular alternative to the VN entropy [14, 18].
-Second, instead of considering that the VN entropy formula is wrong, one may consider that, while the formula is correct, the time evaluation of density matrix $\rho$ is not correctly described by von Neumann-Liouville equation.
+So, the VN entropy is capable of determining the change in (TH) entropy due to a process (the system is in equilibrium before and after the process).
+We also note that this time-invariance of VN entropy is similar to that of the GB entropy.
+The GB entropy of a system is time-invariant since the “shape” of the probability distribution function does not change under the Hamiltonian evolution of the system [7, 19].
+It may be noted that the similarity of the problems faced by the VN and GB entropies is not unexpected, since the VN entropy is the quantum analog of the Gibbs (GB) entropy [20].
+There are many attempts in resolving this issue of time-invariance of the GB entropy under the Hamiltonian evaluation; discussions on them may be found elsewhere [7, 19].
+While trying to resolve the time-invariance problem of the VN entropy, one can think of two major different approaches.
+First, one considers that the VN entropy formula needs modification (or some kind of generalization) so that the entropy from the new formula increases (and saturates in equilibrium) as the system evolves with time.
+The observational entropy is one such popular alternative to the VN entropy [16, 21, 22].
+Second, instead of considering that the VN entropy formula is wrong, one may consider that the time evaluation of the density matrix $\rho$ is not correctly described by the von Neumann-Liouville equation.
 We prefer the second option for the following reasons.
-Like the TH entropy, the VN entropy is a state function - it depends only on $\rho$ and does not depend on any choice or preference of an observer.
-The VN entropy formula correctly captures equilibrium thermodynamics of a system.
+Like the TH entropy, the VN entropy is a state function in the sense that it depends only on the state $\rho$ of the system and does not depend on any choice or preference of an observer.
+The VN entropy formula correctly captures the equilibrium thermodynamics of a system (as mentioned, in equilibrium, $S_{VN}\equiv S_{TH}$).
 In fact, starting with the VN entropy formula, one can derive the expression of the equilibrium Gibbs distribution using the maximum entropy principle and some other conditions [23].
-All these point out that the expression of the VN entropy is a suitable one to study the thermodynamics of a quantum system; main challenge here is to figure out how one the the density matrix evolves from a given initial state to the final one (representing thermal equilibrium).
-In the present study, we leave out any such detail discussion on how the density matrix evolves.
+All these point out that the expression of the VN entropy is a suitable one to study the thermodynamics of a quantum system; main challenge here is to figure out how the density matrix evolves from a given initial state to the final one (representing thermal equilibrium).
+In the present study, we leave out any such detailed discussion on how the density matrix evolves in an irreversible process.
+We just mention here in passing that, although finding an evaluation equation of $\rho$ for an isolated system undergoing an irreversible process can be challenging, such evaluation for an open system is well studied via the Gorini-Kossakowski-Sudarshan-Lindblad equations [24].
+The last thing we may want to note here is that, in some special cases, two approaches mentioned above can actually be equivalent.
+For example, for single coarse-graining, the observational entropy is actually equal to the VN entropy of the coarse-grained density matrix [21, 25].
 Subadditivity of VN entropy:
 The second major objection against the VN entropy is that it obeys the subadditivity property while the TH entropy is additive (for a large uniform system where the interaction energy between its two parts can be ignored).
-The subadditivity property is stated as follows: if a system ($AB$) is in state $\rho_{AB}$, and its parts $A$ and $B$ are respectively in states $\rho_{A}$ and $\rho_{B}$ (which are reduced density matrices of the state $\rho_{AB}$), then $S_{VN}(\rho_{AB})\leq S_{VN}(\rho_{A})+S_{VN}(\rho_{B})$ [20].
+The subadditivity property is stated as follows: if a bipartite system ($AB$) is in state $\rho_{AB}$, and its parts $A$ and $B$ are respectively in states $\rho_{A}$ and $\rho_{B}$ (which are reduced density matrices of the state $\rho_{AB}$), then $S_{VN}(\rho_{AB})\leq S_{VN}(\rho_{A})+S_{VN}(\rho_{B})$ [26].
 This implies that, if the system is in a pure state, its VN entropy will be zero but the sum of the VN entropies of its parts is positive (the sum can be zero in a very special case).
-This contradicts the notion of additivity of the TH entropy.
+This contradicts the notion of the additivity property that the TH entropy has.
 We can circumvent this apparent contradiction in the following way.
-As we noted earlier, the VN entropy is the quantum-analog of the GB entropy.
-The GB entropy is applicable for an open system which is in thermal equilibrium with its bath.
-For such a thermodynamic system, its TH entropy is equal to the TH entropies of its parts.
-Now to verify if the VN entropy has this additivity property, we first note that, the quantum thermalization is defined with respect to a smaller subsystem of a large system [21, 22, 1, 5, 23].
-Here, a part of the system is thermalized while the rest of the system works as its bath.
-Therefore, instead of studying additivity property of the whole quantum system, we should study it for a subsystem.
-It is know that the VN entropy follows a volume law for excited states (especially the states from the middle of its spectrum) [24, 25, 26, 27].
-In fact, for our system (see Sec.
-IV.3), we see that the average VN entropy $\overline{S}_{VN}$ (averaged over all the states from a narrow energy window in the middle of the spectrum) is extensive and grows linearly with subsystem size, as seen in Fig.
+To verify whether the VN entropy has the additivity property, we first note that the quantum thermalization is defined with respect to a subsystem of a (much) larger system [27, 28, 1, 5, 29, 2].
+Essentially one here studies whether the subsystem is thermalized while the rest of the system works as its bath.
+Therefore, instead of studying the additivity property of the whole quantum system, we should study it for a subsystem.
+This is exactly what one does for the classical case where one studies additivity property of the GB entropy (TH entropy) of an open system in thermal equilibrium with its bath.
+For an open system, the additivity implies that the entropies of its parts sum up to the entropy of the system itself (the entropy of the bath does not play any role here).
+With the above understanding, we now check if the VN entropy of a subsystem has the additivity property.
+It is known that the VN entropy follows a volume law for excited states (especially the states from the middle of its spectrum) [30, 31, 32, 33].
+In fact, for our system with $\Delta_{2}=0.5$ (see Sec.
+IV.3), we see that the average VN entropy $\overline{S}^{sb}_{VN}$ (averaged over all the states from a narrow energy window in the middle of the spectrum) is extensive and grows linearly with subsystem size; relevant results can be found in Fig.
 1.
-Figure 1: In the middle of the spectrum, the average VN entropy ($\overline{S}_{VN}$) grows linearly with subsystem size. Here total system size is $N=16$.
+Figure 1: In the middle of the spectrum, the average VN entropy ($\overline{S}^{sb}_{VN}$) grows linearly with subsystem size. Here total system size is $N=16$ and $\Delta_{2}=0.5$.
 We, therefore, infer here that the VN entropy also obeys the additivity property if interpreted appropriately.
@@
 IV Entropy of large quantum systems
-We already saw in Sec.
-II and Sec.
-III.1 that for a thermalized large isolated or open quantum system, the VN entropy is equivalent to the TH entropy (or to an appropriate quantum statistical entropy).
-It is non-trivial to check this hypothesis for a small subsystem of a large quantum system.
-In this section we check, mainly through numerical analysis, whether and how the VN entropy is equivalent to the TH entropy for a subsystem.
-For completeness and more clarity, we first briefly summarize how the VN entropy is equal (or equivalent) to the Th entropy for a thermalized isolated and open system.
-We exclude from our discussion the systems which are not in equilibrium or the systems which do not thermalize.
-Figure 2: The subsystem VN entropy ($S_{VN}$) for individual eigenkets and $\ln(DOS)$ are plotted against the energy of the full system. Here total system size is 16 and subsystem size is 6.
+In Sec.
+III.1, we have seen that the VN entropy of a thermalized isolated or open system is a basis-independent representation of the quantum statistical entropy (which, under certain conditions, is equivalent to the TH entropy).
+There we start with the qBN or qGB entropy and then show that it is simply equal to the VN entropy corresponding to some density matrix (we were not much concerned about any significance or possible interpretation of the density matrix).
+In this section, we go in the opposite direction: we first calculate the VN entropy independently and then verify if that is equal (equivalent) to the TH entropy of the system (subsystem).
+To calculate the VN entropy, we take an appropriate density matrix which describes the system (subsystem) in thermodynamic equilibrium.
+In the following, we discuss the issue of equivalence separately for the isolated, open and subsystem (of a much larger system).
 IV.1 Isolated quantum system
+We consider an isolated quantum system in thermal equilibrium.
 As discussed in Sec.
-II.3, an isolated quantum system will be in a mixed state, for one or multiple reasons.
-Using the principle of equal a priori probabilities, one can show that the VN entropy is equal to the qBN entropy (which in turn is equivalent to the TH entropy).
-Assume that we prepare a system in such a way that its energy lies in the narrow energy shell $(E,E+\Delta E]$, where $\Delta E$ is arbitrary but macroscopically small and microscopically large [5].
-Let {|$\{\ket{n}\}$ be the set of the eigenkets of the system Hamiltonian, and $E_{n}$ be the eigenvalue corresponding to |$\ket{n}$.
-The system will be in any state |$\ket{n}$ with equal probability if $E_{n}\in(E,E+\Delta E]$.
-Accordingly, the thermodynamic entropy ($S_{TH}$), or equivalently, the quantum Boltmann entropy ($S_{qBN}$) is,
+II.3, a large isolated quantum system will be in a mixed state, for one or multiple reasons.
+An isolated system in thermal equilibrium is appropriately described by a microcanonical (mc) ensemble.
+Assume that we prepare the system in such a way that its energy lies in the narrow energy shell $(E,E+\Delta E]$, where $\Delta E$ is arbitrary but macroscopically small and microscopically large [5].
+Let {|$\{\ket{n}\}$ be the set of eigenkets of the Hamiltonian of the system, and $E_{n}$ be the eigenvalue corresponding to |$\ket{n}$.
+The density matrix which describes the isolated system is given by,
+$\rho_{mc}=\sum_{\begin{subarray}{c}n\\ E_{n}\in\left(E,E+\Delta E\right]\end{subarray}}p_{n}\ket{n}\bra{n},$ (8)
+where $p_{n}$ is the probability to find the system in state |$\ket{n}$.
+Since $\Delta E$ is macroscopically small, all the accessible microstates (here, the eigenkets) will essentially have almost the same energy.
+Application of the principle of equal a priori probabilities suggests that $p_{n}=\frac{1}{D}$ (see aslo Sec.
+II.3).
+As $S_{VN}=-k_{B}\text{Tr}(\rho_{mc}\ln\rho_{mc})$, we need the eigenvalues of $\rho_{mc}$ to efficiently compute $S_{VN}$.
+It is clear from Eq.
+8 that |$\ket{n}$ is an eigenket of $\rho_{mc}$ with eigenvalue $\frac{1}{D}$, i.e., $\rho_{mc}\ket{n}=\frac{1}{D}\ket{n}$ for any energy eigenket from the shell.
+This implies that
+$S_{VN}=-k_{B}\text{Tr}(\rho_{mc}\ln\rho_{mc})=k_{B}\ln D.$ (9)
+If we now directly calculate the TH entropy (or equivalently, the qBN entropy), following our discussion in Sec. II.3, it would be
 $S_{TH}\equiv S_{qBN}=k_{B}\ln D,$ (10)
@@
 For a thermodynamically large system, the dimension of this largest subspace, which corresponds to the thermal equilibrium, is essentially equal to the dimension of the shell ($D$).
-Now to calculate the corresponding von Neumann entropy ($S_{VN}$) of the isolated quantum system, we note that the system is in a mixed state, represented by the following density matrix,
-$\rho=\sum_{\begin{subarray}{c}n\\ E_{n}\in\left(E,E+\Delta E\right]\end{subarray}}p_{n}\ket{n}\bra{n},$ (9)
-where $p_{n}$ is the probability to find the system in state |$\ket{n}$.
-Since $\Delta E$ is macroscopically small, all the accessible microstates (here, eigenkets) will essentially have almost the same energy.
-Application of the principle of equal a priori probabilities suggests that $p_{n}=\frac{1}{D}$.
-As $S_{VN}=-k_{B}\text{Tr}(\rho\ln\rho)$, we need the eigenvalues of $\rho$ to efficiently compute $S_{VN}$.
-It is clear from Eq.
-9 that |$\ket{n}$ is an eigenket of $\rho$ with eigenvalue $\frac{1}{D}$, i.e., $\rho\ket{n}=\frac{1}{D}\ket{n}$ for any energy eigenket from the shell.
-This implies that
-$S_{VN}=-k_{B}\text{Tr}(\rho\ln\rho)=k_{B}\ln D.$ (10)
 We see from Eqs.
 9 and 10 that, for an isolated quantum system, the VN entropy is the same as the TH entropy ($S_{VN}=S_{qBN}\equiv S_{TH}$), provided that the system is large and in thermal equilibrium.
-The detail discussion on the conditions under which this equivalence works can be found in Sec.
+The detailed discussion on the conditions under which this equivalence works can be found in Sec.
 II.2.
 IV.2 Weakly interacting open quantum system
-Here we consider an open system which is in contact with the thermal bath.
-Although the system exchanges heat with the bath, we assume that the interaction with the bath is so weak that the energy spectrum of the system is more or less unaffected by the interaction, and the dynamics of the quantum system can still be described (at-least effectively) by its Hamiltonian.
-As before, let $\{E_{n}\}$ be the set of eigenvalues of the system Hamiltonian with |$\ket{n}$ being the eigenket corresponding to eigenvalue $E_{n}$.
-In the spirit of Gibbs distribution for classical system, we can define the probability to find the system with energy $E_{n}$ as $p_{n}=\frac{1}{Q}e^{-\beta E_{n}}$, where $Q$ is the partition function, $Q=\sum_{n}e^{-\beta E_{n}}$.
-Here $\beta$ denotes the inverse temperature.
-We can now get the TH entropy, quantified by the quantum Gibbs entropy, as,
-$S_{TH}\equiv S_{qGB}=-k_{B}\sum_{n}p_{n}\ln{p_{n}}.$ ) . (11)
-Figure 3: The average VN entropy ($\overline{S}_{VN}$) for subsystem is plotted against $\ln(DOS)$ across the energy spectrum (separately for the left and the right halves of the spectrum).
-Our system size is 16 and subsystem size is 6.
-To find the VN entropy of the open system, one takes the following density operator: $\rho=\frac{e^{-\beta H}}{Tr(e^{-\beta H})}$, where the inverse temperature can be fixed by solving the equation $Tr(He^{-\beta H})=U$ with $U$ being the (given) internal or average energy.
-In fact, for overwhelmingly large number of the pure states (of the system-bath combination), it can be quite generally shown that, a weakly interacting open quantum system is very well described by a (reduced) density matrix which is equivalent to the known Gibbsian distribution [29, 30].
-As a result, the VN entropy of the open system is,
-$S_{VN}=-k_{B}\text{Tr}(\rho\ln\rho)=-k_{B}\sum_{n}p_{n}\ln{p_{n}},$ (12)
-where the expression of $p_{n}$ is the same as the one appears in the formula for $S_{qGB}$ in Eq.
-11.
+Here we consider an open system that is in contact with the thermal bath.
+Although the system exchanges heat with the bath, we assume that the interaction with the bath is so weak that the energy spectrum of the system is more or less unaffected by the interaction, and the dynamics of the quantum system can still be described (at-least effectively) by its Hamiltonian $H$.
+As discussed in Sec.
+II.3, the density matrix that appropriately describes the open system is $\rho_{\beta}=\frac{e^{-\beta H}}{Tr(e^{-\beta H})}$, where the inverse temperature can be fixed by solving the equation $Tr(He^{-\beta H})=U$ with $U$ being the (given) internal or average energy.
+In fact, if the system-bath combination is described by a microcanonical ensemble or for the overwhelmingly large number of pure states (of the system-bath combination), it can be quite generally shown that an open quantum system is very well described by a (reduced) density matrix which is equivalent to the known Gibbsian distribution [11, 10].
+Accordingly, the VN entropy of the open system is,
+$S_{VN}=-k_{B}\text{Tr}(\rho_{\beta}\ln\rho_{\beta})=-k_{B}\sum_{n}p_{n}\ln{p_{n}},$ (11)
+where $p_{n}=\frac{1}{Q}e^{-\beta E_{n}}$ with $Q$ being the partition function, $Q=\sum_{n}e^{-\beta E_{n}}$. Here, $\{E_{n}\}$ is the set of eigenvalues of the system Hamiltonian $H$.
+Figure 2: The subsystem VN entropy ($S^{sb}_{VN}$) for the individual eigenkets and $\ln(DOS)$ are plotted against the energy of the full system. Here total system size is 16 and subsystem size is 6.
+As discussed in Sec. II, for the open quantum system, the TH entropy (or equivalently, the qGB entropy) is given by:
+$S_{TH}\equiv S_{qGB}=-k_{B}\sum_{n}p_{n}\ln{p_{n}},$ (12)
+where the expression of $p_{n}$ is the same as the one appears in Eq. 11.
+Figure 3: The average VN entropy ($\overline{S}^{sb}_{VN}$) of subsystem is plotted against $\ln(DOS)$ across the energy spectrum (separately for the left and the right halves of the spectrum).
+Here, the system (subsystem) size is 16 (6) and $\Delta_{2}=0.5$.
 We now see from Eqs.
-11 and 12 that the the VN entropy is the same as the TH entropy ($S_{VN}=S_{qBN}\equiv S_{TH}$) for a weakly interacting open quantum system in thermal equilibrium.
-The detail discussion on the conditions under which this equivalence works can be found in Sec.
+11 and 12 that the VN entropy is the same as the TH entropy ($S_{VN}=S_{qBN}\equiv S_{TH}$) for a weakly interacting open quantum system in thermal equilibrium.
+The detailed discussion on the conditions under which this equivalence works can be found in Sec.
 II.2.
 IV.3 Subsystem of a quantum system
-As mentioned in the beginning of this section (Sec.
-IV), it is non-trivial to verify if the VN entropy is equivalent to the TH entropy for a subsystem (of a much larger quantum system in thermal equilibrium).
-We analyze this possibility here.
-We consider a large composite quantum system $AB$ in a state described by the density matrix $\rho$.
-We also assume that the subsystem $A$ is macroscopically large but much smaller compare to its counterpart $B$.
-We here show, mainly by the numerical analysis, that the VN entropy of the subsystem $A$ is proportional to its TH entropy.
-To first calculate the TH entropy of a subsystem, we again assume that the total system is prepared with an energy lying in $(E,E+\Delta E]$.
-As discussed earlier, the system will be in a mixed state $\rho$, as appear in Eq.
-9.
-The TH entropy (qBN entropy) of the full system is $S_{TH}\equiv S_{qBN}=k_{B}\ln D$, as in Eq.
-8, where the $D$ is the number of eigenvalues of the system lying in the aforementioned energy shell.
+We separately discuss here the question of equivalence between the subsystem entropies when the full system is in a mixed state (described by a microcanonical ensemble) or in a pure state (both typical and atypical).
+If the full system is in a mixed state $\rho_{mc}$, described by a microcanonical (mc) ensemble, then the subsystem will be in a Gibbsian distribution $\rho_{\beta}$ [11, 10].
+If $\rho_{sb}$ is the reduced density matrix for the subsystem, i.e., $\rho_{sb}=\text{Tr}_{B}(\rho_{mc})$, then $\rho_{sb}\approx\rho_{\beta}=\frac{e^{-\beta H}}{Q}$ with $Q=\text{Tr}_{sb}(e^{-\beta H})$.
+It may be noted that the first trace (Tr) operation here is over the degrees of freedom of the bath, B (complement of the subsystem in the full system), and the second trace operation is over the degrees of freedom of the subsystem (sb).
+We also note that, here, $H$ is the Hamiltonian of the subsystem if the interaction between the subsystem and the bath can be ignored, else it is the effective Hamiltonian of the subsystem, determined by $H=-\frac{1}{\beta}\ln(Q\text{Tr}_{B}(\rho_{mc}))$.
+This equation can be solved self-consistently: initially, we calculate $Q$ from the Hamiltonian of the subsystem (without interaction with the bath) and get the first order approximation of the effective $H$.
+Next we use this effective $H$ in $Q$ to find the next order approximation of the effective $H$.
+We continue this till we get a self-consistent result.
+Fortunately, we do not need to calculate this effective $H$ in the numerical calculations that follow.
+If the full system is in a pure state, for the overwhelming majority of pure states (of the total system), the subsystem can be described by a Gabssian distribution $\rho_{\beta}$ [11, 10].
+This important result, known as canonical typicality, leaves open the possibility of pure states (of the full system) for which a subsystem is not described by a Gibbsian distribution or for which a subsystem does not thermalize.
+We note that there are systems whose eigenstates (from a narrow energy window) are such that the reduced density matrices calculated from them are quite different to each other.
+These systems do not follow the eigenstate thermalization hypothesis or ETH [27, 28, 5].
+Above important results immediately imply the following: if the full system is described by a microcanonical ensemble, the VN entropy of a subsystem is equivalent to its TH entropy (qGB entropy).
+If the full system is in a pure state, for the overwhelming cases (i.e.
+for the typical states), the VN entropy of a subsystem will be equivalent to its TH entropy (qGB entropy).
+But for systems for which the ETH does not hold, this equivalence will not work in general (corresponding pure states are atypical).
+In the following we take a spin system and numerically demonstrate the issue of equivalence between the VN entropy and TH entropy of a subsystem in the different cases mentioned above.
+For our numerical results, we take the following one-dimensional (with open boundary condition) Heisenberg spin-1/2 chain with the next-nearest neighbor interactions:
+$H=\sum_{i=1}^{N-1}\vec{S}_{i}\cdot\vec{S}_{i+1}+\Delta_{2}\sum_{i=1}^{N-2}S^{z}_{i}S^{z}_{i+2}.$ (13)
+In our calculations, we take the size of the system $N=16$. This model system is integrable when $\Delta_{2}=0$ and is non-integrable (hence thermalizes) when $\Delta_{2}>0$ [35].
+To numerically calculate the VN entropy of the subsystem, we first find $\rho_{sb}$ from the state ($\rho$) of the full system: $\rho_{sb}=\text{Tr}_{B}(\rho)$.
+We then use the formula Eq.
+3 to calculate the VN entropy of the subsystem.
+To numerically calculate the corresponding TH entropy, consider that the full system is prepared with an energy lying in $(E,E+\Delta E]$.
+As discussed earlier, the system will be in a mixed state $\rho_{mc}$, as appear in Eq.
+8.
+The TH (or qBN) entropy of the full system is $S_{TH}\equiv S_{qBN}=k_{B}\ln D$, as in Eq.
+10, where $D$ is the number of eigenvalues of the system lying in the aforementioned energy shell.
 Now since the TH entropy is additive, we can assume that the TH entropy of the subsystem (denoted as $S^{sb}_{TH}$ ) is proportion to $k_{B}\ln D$ or to the $k_{B}\ln(DOS)$, where $DOS$ is the density-of-states of the full system at the energy scale $E$.
 Accordingly,
-$S^{sb}_{TH}\propto k_{B}\ln(DOS),$ (13)
+$S^{sb}_{TH}\equiv S^{sb}_{qBN}\propto k_{B}\ln(DOS),$ (14)
 where the proportionality constant depends on the size of the subsystem in comparison to the size of the full system.
-Now to calculate the VN entropy of the subsystem, we first need to calculate the reduced density matrix of the subsystem $\rho_{A}=\text{Tr}_{B}(\rho)$, where $\rho$ is the density matrix of the full system, as given in Eq.
-9 (with $p_{n}=\frac{1}{D}$).
-Denoting the reduced density matrix corresponding to the eigenket |$\ket{n}$ by $\rho_{A}^{(n)}$, where $\rho_{A}^{(n)}=\text{Tr}_{B}(\ket{n}\bra{n})$, we have
-$\rho_{A}=\frac{1}{D}\sum_{n}\rho_{A}^{(n)}.$ (14)
-Since the subsystem is in the mixed state $\rho_{A}$, we now can calculate the VN entropy of the subsystem directly from $\rho_{A}$, i.e,
-$S^{sb}_{VN}=-k_{B}\text{Tr}(\rho_{A}\ln\rho_{A}),$ (15)
-where $\rho_{A}$ is given by Eq. 14. For a system which thermalizes (or obeys the eigenstate thermalization hypothesis [21, 22]), we can also have
-$S^{sb}_{VN}=-k_{B}\text{Tr}(\rho_{A}\ln\rho_{A})\approx-\frac{k_{B}}{D}\sum_{n}\text{Tr}(\rho_{A}^{(n)}\ln\rho_{A}^{(n)}).$ (16)
-This approximate equality works since, as per the hypothesis, the expectation value of a local operator (say, $O$) is almost the same for different eigenkets from a narrow energy window, i.e., $\text{Tr}(\rho_{A}^{(n)}O)\approx\text{Tr}(\rho_{A}^{(m)}O)$.
-This is possible if the set of eigenvalues of the reduced density matrix $\rho_{A}^{(n)}$ is almost the same for different $n$.
-This implies that $\text{Tr}(\rho_{A}^{(n)}\ln\rho_{A}^{(n)})\approx\text{Tr}(\rho_{A}^{(m)}\ln\rho_{A}^{(m)})$, which guarantees the approximate equality in Eq.
+Now, for different cases, we calculate the VN entropy and the TH entropy of a subsystem to study their equivalence.
+First, we consider the case when the full system is in an eigenket |$\ket{n}$ of the Hamiltonian of Eq.
+13.
+In this case the subsystem will be represented by the reduced density matrix $\rho_{sb}\equiv\rho_{sb}^{(n)}=\text{Tr}_{B}(\ket{n}\bra{n})$.
+Now using Eq.
+3, we can calculate the VN entropy of the subsystem when the full system is in the eigenket |$\ket{n}$:
+$S^{sb}_{VN}=-k_{B}\text{Tr}(\rho_{sb}\ln\rho_{sb}).$ (15)
+The values of $S^{sb}_{VN}$ for different eigenkets across the spectrum can be found in Fig.
+2 (for $\Delta_{2}=0.0,0.2$ and $0.5$).
+We also show the corresponding logarithm of the density-of-states (DOS) for the full system.
+As mentioned earlier, when the system does not thermalize ($\Delta_{2}=0.0$), the individual reduced density matrices are quite different from each other (even within a narrow energy shell) and accordingly, $S^{sb}_{VN}$ takes quite different values as one goes from one eigenket to another within the energy shell.
+In this case the VN entropy of the subsystem is not, in general, proportional to its qBN entropy, i.e., to $\ln(DOS)$.
+On the other hand, when the system thermalizes (or the ETH holds), $S^{sb}_{VN}$ takes very close values as one goes from one eigenket to another within an energy shell (see, for example, results for $\Delta_{2}=0.5$).
+In Fig.
+3, we plot the average VN entropy ($\overline{S}^{sb}_{VN}$) of the subsystem against the logarithm of DOS (individually for left and right halves of the spectrum).
+We take the average of $S^{sb}_{VN}$ over a small energy shell to get better numerical results.
+The linearity of the plots shows that the VN entropy of a subsystem is equivalent to its TH (or qBN) entropy, i.e., $S^{sb}_{VN}\approx\overline{S}^{sb}_{VN}\equiv S^{sb}_{qBN}$.
+Next, we consider the case when the full system is in a mixed state ($\rho_{mc}$), as appears Eq. 8 (with $p_{n}=\frac{1}{D}$). In this case, the subsystem will be in the following mixed state,
+$\rho_{sb}=\frac{1}{D}\sum_{n}\rho_{sb}^{(n)},$ (16)
+where the reduced density matrix $\rho_{sb}^{(n)}=\text{Tr}_{B}(\ket{n}\bra{n})$ is calculated for a single eigenket |$\ket{n}$ of the full system.
+Here the sum is over the all eigenkets from a small (fixed) energy shell.
+There are $D$ eigenkets in the shell.
+Using Eq.
+15, we now can calculate the VN entropy of the subsystem directly from $\rho_{sb}$ of Eq.
 16.
-If we write $\overline{S}_{VN}=-\frac{k_{B}}{D}\sum_{n}\text{Tr}(\rho_{A}^{(n)}\ln\rho_{A}^{(n)})$, then $S^{sb}_{VN}\approx\overline{S}_{VN}$ for a thermalized system.
-We note that, if the system does not thermalize (or is integrable), then the VN entropy for the individual eigenkets ($-k_{B}\text{Tr}(\rho_{A}^{(n)}\ln\rho_{A}^{(n)})$) can take widely different values.
-This in fact can be seen in Fig.
-2, where the subsystem VN entropy for different eigenkets are plotted against the energy across the spectrum.
-We here mention that, for all our numerical results, we take the following one-dimensional (with open boundary condition) Heisenberg spin-1/2 chain with the next-nearest neighbor interactions:
-$H=J\left(\sum_{i=1}^{N-1}\vec{S}_{i}\cdot\vec{S}_{i+1}+\Delta_{2}\sum_{i=1}^{N-2}S^{z}_{i}S^{z}_{i+2}\right).$ (17)
-In our calculations, we take $J=1$ and the size of the system $N=16$. This model system is integrable when $\Delta_{2}=0$ and is non-integrable (hence thermalizes) when $\Delta_{2}>0$ [31].
-To make more sense of the results in Fig.
-2, we plot the subsystem’s average VN entropy ($\overline{S}_{VN}$) against $\ln(DOS)$ (separately for the left and right halves of the spectrum).
-The results are shown in Fig.
+The value of $S^{sb}_{VN}$ across the energy spectrum can be found in Fig.
+4.
+Figure 4: The subsystem VN entropy ($S^{sb}_{VN}$) when the full system is in the microcanonical density matrix $\rho_{mc}$. Here, the system (subsystem) size is 16 (6).
+The plot of $S^{sb}_{VN}$ against $\ln(DOS)$ shows linearity, similar to what we find in Fig.
 3.
-We find that, for a thermalized system ($\Delta_{2}>0$), its subsystem’s average VN entropy ($\overline{S}_{VN}$) is proportional to $\ln(DOS)$ with a proportionality constant approximately equal to $\frac{N_{sb}}{N}$ (in our calculations, the subsystem size $N_{sb}=6$ and the full system size $N=16$).
-This proportionality constant takes a different value (a smaller one) if the system is not thermalized.
-We take $k_{B}=1$ for all our numerical calculations of entropy.
-Thus our numerical results suggest that,
-$S^{sb}_{VN}\approx\overline{S}_{VN}\propto\ln(DOS).$ (18)
-From Eqs. 13 and 18, we conclude that, like for isolated and open systems, for a subsystem also the VN entropy is equivalent to the TH entropy (qBN entropy), i.e., $S^{sb}_{VN}\equiv S^{sb}_{TH}$.
+We note that, as expected, we get the said linear plot even when the system is in integrable regime ($\Delta_{2}=0.0$).
 It may be worth mentioning here that, the result that the subsystem’s average VN entropy is proportional to system’s $\ln(DOS)$, is earlier reported in [36] in a different context.
-We also note that, elsewhere [33], the thermodynamic entropy of a subsystem is defined as what we call here $S^{sb}_{VN}$.
-As discussed in this work, thermodynamic (TH) entropy is appropriately calculated as the logarithm of $DOS$ (since, $S_{TH}\equiv S_{qBN}$ under some general conditions).
+We also note that, elsewhere [37], the TH entropy of a subsystem is defined as what we call here the VN entropy of a subsystem ($S^{sb}_{VN}$) when the full system is in $\rho_{mc}$.
+As discussed in this work, the TH entropy of a subsystem is appropriately calculated as the logarithm of $DOS$ (since, $S^{sb}_{TH}\equiv S^{sb}_{qBN}\propto\ln(DOS)$ under some general conditions).
 V Conclusion
@@
 In reviewing the major objections about the VN entropy, we explain that the concerns about the time-invariance and subadditivity properties can be addressed convincingly.
-We show in this study that, for large thermalized isolated or open system, the VN entropy is equivalent to the TH entropy.
-Next, through numerical analysis, we also show that the VN entropy of a subsystem of a large thermalized system is equivalent to its TH entropy (qBN entropy).
+We find that, when a large isolated system is in a mixed state (described by microcanonical ensemble), its VN entropy is equivalent to its TH entropy.
+When the system-bath combination is in a mixed state, or in a typical pure state, the VN entropy of the (open) system is equivalent to its TH entropy.
+This result also holds for a subsystem (of a much larger system) but the analytical demonstration is not straightforward.
+Through numerical analysis, we show that the VN entropy of a subsystem is equivalent to its TH (qBN) entropy when the full system (subsystem-bath combination) is described by a microcanonical ensemble or when the full system is in a typical pure state.
+In this work we also numerically demonstrate that, for the states that are atypical (or for which the ETH does not hold), the VN entropy and the TH entropy of a subsystem are not the same.
+Acknowledgements.
+SS thanks S. Aravinda for useful discussions.

### v3

@@
-von Neumann entropy and quantum version of thermodynamic entropy
+Statistical entropy of quantum systems
 Abstract
-The debate whether the von Neumann (VN) entropy is a suitable quantum version of the thermodynamic (TH) entropy has a long history.
-In this regard, we briefly review some of the main reservations about the VN entropy and explain that the objections about its time-invariance and subadditivity properties can be addressed convincingly.
-In a broader context, we here analyze in detail whether and when the VN entropy is the same or equivalent to the TH entropy.
-We find that the VN entropy of a large isolated system is equivalent to its TH entropy if the system is in a mixed state described by a microcanonical (MC) ensemble.
-This equivalence also holds for open system and subsystem if the full system (i.e.
-system-bath or subsystem-bath combination) is in a mixed state described by an MC ensemble or in a typical pure state (or if the Eigenstate Thermalization Hypothesis, ETH, holds for the full system).
-Unlike isolated and open system, demonstrating the equivalence for a subsystem is not straightforward.
-We take a one-dimensional spin-1/2 model to numerically show that, if the full system is in a pure state, then the VN entropy of a subsystem is equivalent to its TH entropy if the ETH holds for the full system.
-We also demonstrate that this equivalence always works (irrespective of whether the ETH holds) if the full system is in a mixed state describe by an MC ensemble.
-preprint: APS/123-QED
-I Introduction
+Let $D_{1}$ and $D_{2}$ be the Hilbert space dimensions of two subsystems of a quantum system of total Hilbert space dimension $D=D_{1}D_{2}$.
+In the thermodynamic limit (with $1\ll D_{1}\ll D_{2}$), we know from the works of Page and Sen that the average von Neumann (VN) entropy of the first subsystem is $\mathbb{E}({S}^{sb}_{VN})=\ln(D_{1})+O(D_{1}/D_{2})$ if the full system is in a random pure state.
+Here, it is argued that this result can be strengthened for a thermalized quantum system.
+Consider the subspace $\mathcal{H}_{E}$ of the total Hilbert space corresponding to a narrow shell around the energy $E$.
+We find that the result of Page and Sen holds for each of these subspaces, that is, the VN entropy, when averaged over the states in $\mathcal{H}_{E}$, is given by $\overline{S}^{sb}_{VN}\approx\ln\widetilde{d}_{1}$, where $\widetilde{d}_{1}$ represents the dimension of the effective Hilbert space of the first subsystem relevant to $\mathcal{H}_{E}$.
+If $d_{E}=\dim{(\mathcal{H}_{E})}$, we estimate that $\widetilde{d}_{1}=D_{1}^{\gamma}$, where $\gamma=\ln(d_{E})/\ln(D)$.
+This finding has significant implications, as it suggests an equivalence between the VN entropy and the thermodynamic (TH) entropy of a subsystem within a much larger thermalized quantum system.
+For completeness, we also discuss in this work the issue of equivalence between the VN entropy and TH entropy for isolated system (as a whole) and open system.
+For numerical demonstration of our important results, we here consider a one-dimensional spin-1/2 chain with next-nearest neighbor interactions.
+Keywords: Entropy of quantum systems, von Neumann entropy, Thermodynamic entropy
+1 Introduction
 Since the development of quantum mechanics, there has been an enormous progress in our understanding of how the quantum many-body systems behave.
@@
 The thermodynamic (TH) entropy is a very important state function which helps us understand how a system approaches equilibrium besides playing a crucial role in studying equilibrium properties of a system [16].
-The usefulness of the TH entropy naturally calls for its appropriate quantum counterpart in QSM.
-The TH entropy, as defined by Clausius in the context of second law of thermodynamics, is a state function which takes the maximum value for a system in the equilibrium under fixed macroscopic conditions [4, 5].
-For an isolated system, this entropy never decreases when the system undergoes some process.
-Thus any entropy defined in the similar spirit for a quantum system must be a state function and should increase (or remain constant) when the system undergoes a process.
-Boltzmann, and later Gibbs, showed us how the TH entropy can be calculated from the microscopic details using statistical approaches [4, 6].
-For an isolated system, Boltzmann gave a formula (actually Planck first wrote down the explicit expression for the Boltzmann entropy) for the TH entropy in terms of the volume of the relevant phase space (or the number of associated microstates).
-For the open system, Gibbs on the other hand, using his ensemble theory, provided the formula for the TH entropy in terms of an appropriate probability distribution function in phase space.
-Although the quantum versions of these statistical entropies for the isolated and open systems in equilibrium are somewhat intuitively clear [7, 5], one needs convincing justifications, which we discuss in Sec.
-II.
-The question, however, remains on the quantum version of the TH entropy for a part (subsystem) of a quantum system.
-In this work, we will discuss this issue in detail.
-As noted, for a quantum system in thermal equilibrium, writing down the quantum version of the Gibbs (GB) entropy is somewhat easy (see, Secs.
-II.3 and III.1).
-But when a quantum system is in a generic mixed state and not in equilibrium, what would be the value of entropy?
-From thermodynamical considerations, von Neumann (VN) gave a formula for the entropy of a quantum ensemble (representing a given mixed state) [8].
-This entropy, known as VN entropy, attracted a great deal of attention and a lot of criticism in regards to its connection to the TH entropy.
-We show in this work that, for isolated and open quantum system or for a subsystem of a (much) larger quantum system, the VN entropy is equivalent to the TH entropy if the quantum system (or subsystem) is large and in thermal equilibrium.
-We largely exclude from our discussions the quantum systems that do not thermalize or systems that are not in equilibrium.
-We here, though, deliberate on two main criticisms on the VN entropy, related to the time-invariance and subadditivity, and argue that these criticisms can be avoided or addressed convincingly.
-We outline here the logical steps we follow in this work while establishing the said equivalence between the VN and TH entropies and finding the conditions under which the equivalence works.
-First, we identify under what conditions the TH entropy is equivalent to the statistical entropy (either Boltzmann’s or Gibbs’ version), then reinterpret the statistical entropy in terms of the quantum variables (in Boltzmann’s case: the dimension of relevant subspace in place of the volume of accessible phase-space, and in Gibbs’ case: the energy eigenvalues in place of the continuous classical energy).
-In the last step of establishing the equivalence, we check if the VN entropy of an appropriate density matrix is equivalent to the quantum version of the statistical entropy (i.e.
-the reinterpreted statistical entropy in terms of quantum variables).
-To better understand the equivalence, one should also check if the VN entropy inherits the important properties of the equilibrium TH entropy.
-Accordingly, wherever appropriate, we briefly deliberate if the VN entropy is a state function, has the additivity property and whether it increases for isolated system due to an irreversible process.
-We organize our paper in the following way.
-In the next section (Sec.
-II), we review the thermodynamic and statistical entropies, and discuss the condition of equivalence between these entropies.
-This is followed by a discussion on the suitable quantum versions of the statistical entropies.
+The Boltzmann (BN) entropy and Gibbs (GB) entropy are two well-established statistical measures that represent the thermodynamic (TH) entropy of classical systems.
+The success of these statistical entropies inspires the search for similar statistical quantities that can accurately represent the TH entropy of quantum systems.
+In this regard, a key question is whether the von Neumann (VN) entropy is a suitable candidate for this role.
+Against this backdrop, here we explore some key challenges and open questions in the field of QSM.
+Our primary contributions are as follows.
+We argue here that the landmark result of Page [22] and Sen [30] can be strengthened for a thermalized system.
+Their result roughly shows that the von Neumann (VN) entropy of a subsystem, when the full system is in a random pure state, is equal to the logarithm of the dimension of the subsystem’s Hilbert space.
+For a thermalized system, we show that a result equivalent to that of Page and Sen holds for any subspace of the full Hilbert space associated with a narrow shell around some energy $E$.
+We discuss the above result in a broader context.
+As mentioned in the beginning, it is of fundamental interest to know whether the VN entropy is a suitable statistical measure of the TH entropy of a quantum system.
+The above-mentioned result answers an important part of this general question: it is argued that this result implies an equivalence between the VN entropy and the TH entropy for a subsystem when the full (thermalized) system is in a (random) pure state associated with a narrow energy shell.
+This question of equivalence is also addressed using the notion of canonical typicality [26, 11].
+We also address the question of equivalence between the VN entropy and the TH entropy of a subsystem when the full system is in a mixed state.
+We emphasize that understanding the equivalence between the VN entropy and the TH entropy of a subsystem is crucial in the context of quantum thermalization, as thermalization is essentially defined with respect to a subsystem (or a local operator) [19, 5, 33].
+For completeness and better understanding, the same question of equivalence is also addressed for open system (with very weak interaction with the environment) and isolated system (as a whole).
+For demonstration of the nontrivial results, we also perform numerical analysis on a spin-1/2 chain with next-nearest neighbor interactions.
+Since understanding the equivalence between the VN entropy and the TH entropy (in any of the cases mentioned above) requires evaluating both the entropies separately, we first, in Sec.
+2, deliberate on how one can quantify the TH entropy of a quantum system.
+We also provide a detailed discussion on the conditions under which this quantification holds.
+Next, in Sec.
+3, we have a general discussion on the VN entropy and address two major objections, related to time invariance and subadditivity properties, when compared to the TH entropy in a quantum system.
 In Sec.
-III, we discuss von Neumann entropy and some reservations about the entropy when it comes to representing the thermodynamic entropy.
-Next, in Sec.
-IV, we check for three different cases (isolated system, open system and subsystem) if the von Neumann entropy is equivalent to the thermodynamic entropy.
+4, we discuss the issue of equivalence between the VN entropy and the TH entropy separately for three cases: isolated system, open system and subsystem of a larger system.
 We conclude our work in Sec.
-V.
-II Thermodynamic and statistical entropies
-In this section we first briefly review the contexts and definitions of the thermodynamic (TH) entropy and the statistical (ST) entropies (both Boltzmann’s and Gibbs’ versions).
+5.
+2 Thermodynamic and statistical entropies
+In this section we first briefly review the thermodynamic (TH) entropy and the statistical (ST) entropies (both Boltzmann’s and Gibbs’ versions).
 Next we try to understand under what conditions these entropies (TH and ST) are equivalent.
-At the end this section, we discuss the reinterpretations of the statistical entropies in terms of the quantum variables.
-II.1 Entropy: from Clausius to Boltzmann and Gibbs
+At the end of this section, we discuss the reinterpretations of these statistical entropies in terms of the quantum variables.
+2.1 Entropy: from Clausius to Boltzmann and Gibbs
 The Clausius theorem or the Clausius inequality states that [16], for any cyclic transformation or thermodynamic process, $\oint\frac{d\bar{}\hskip 0.70004ptq}{T}\leq 0$, where $d\bar{}\hskip 1.00006ptq$ is the heat increment received by the system at temperature $T$.
@@
 For a gaseous system, this number is proportional to the phase-space volume available to the system in the given macrostate.
-Here the statistical entropy is defined as $S_{BN}=k_{B}\ln{\Omega}$ ) with $k_{B}$ being the Boltzmann (BN) constant [4, 9, 7].
+Here the statistical entropy is defined as $S_{BN}=\ln{\Omega}$ (in this paper we set the Boltzmann constant $k_{B}=1$) [16, 17, 12].
 By dividing the accessible phase-space as per the values of the well-defined macrovariables, one can in-fact calculate the BN entropy for each such division [12, 19].
@@
 In the following subsection we discuss under what conditions the TH entropy and the statistical entropy (BN or GB entropy) are equivalent.
-II.2 Conditions of equivalence between thermodynamic and statistical entropies
+2.2 Conditions of equivalence between thermodynamic and statistical entropies
 To understand the conditions under which the equivalence works, we outline here a method to establish such equivalence.
 A common way of showing the equivalence between the TH and BN entropies of a system with a given energy is to bring the system with energy $E_{1}$ in contact with another system with energy $E_{2}$ [23].
-Before the contact, the systems were in thermal equilibrium and after the contact they will again come to a common equilibrium via exchange of heat.
+Before the contact, the systems were individually in thermal equilibrium and after the contact they will again come to a common equilibrium via exchange of heat.
 Although the systems exchange heat, the total energy ($E=E_{1}+E_{2}$) of the composite system remains constant; here one ignores the interaction energy between the systems.
@@
 […]
-II.3 Quantum version of statistical entropy
+2.3 Reinterpreting statistical entropies for quantum systems
 Intuitively it is not very difficult to guess the quantum version of a statistical entropy.
@@
 Once we have agreement on the quantum version of the principle of equal a priori probabilities, it is not difficult to come up with quantum versions of the statistical entropies.
-To get the quantum version of the BN (qBN) entropy, we prepare a system in such a way that its energy lies in the narrow energy shell $(E,E+\Delta E]$, where $\Delta E$ is arbitrary but macroscopically small and microscopically large [5].
+To get the quantum version of the BN entropy (qBN entropy), we prepare a system in such a way that its energy lies in the narrow energy shell $(E-\frac{\Delta E}{2},E+\frac{\Delta E}{2}]$, where $\Delta E$ is arbitrary but macroscopically small and microscopically large [19].
 Let $\{E_{n}\}$ be the set of eigenvalues of the system Hamiltonian and {|$\{\ket{n}\}$ be the corresponding set of eigenkets.
-As per the premise, the system will be in any state |$\ket{n}$ with equal probability if $E_{n}\in(E,E+\Delta E]$.
+As per the premise, the system will be in any state |$\ket{n}$ with equal probability if $E_{n}\in(E-\frac{\Delta E}{2},E+\frac{\Delta E}{2}]$.
 Subsequently, guided by our understanding of the BN entropy, we can write its quantum version ($S_{qBN}$) in the following way:
-$S_{BN}\to S_{qBN}=k_{B}\ln D,$ (1)
-where $D$ is the number of eigenvalues within the shell.
-It may be noted here that one can subdivide the energy shell into different macrostates (subspaces), with one of them having the largest number of microststes (the largest dimension) [7, 5].
-For a thermodynamically large system, the dimension of this largest subspace, which corresponds to the thermal equilibrium, is essentially equal to the dimension of the shell ($D$).
+$S_{BN}\to S_{qBN}=\ln d_{E},$ (1)
+where $d_{E}$ is the number of eigenkets (microstates) associated with the shell.
+It may be noted here that one can subdivide the energy shell into different macrostates (subspaces), with one of them having the largest number of microststes [12, 19].
+For a thermodynamically large system, the dimension of this largest subspace, which corresponds to the thermal equilibrium, is essentially equal to $d_{E}$.
 To get the quantum version of the Gibbs (qGB) entropy, in place of continuous classical energy, we use energy eigenvalues of the system’s Hamiltonian to calculate the partition function.
@@
 Here $\beta$ denotes the inverse temperature.
-Subsequently, we can write the quantum version of the GB entropy in the following way:
-$S_{GB}\to S_{qGB}=-k_{B}\sum_{n}p_{n}\ln{p_{n}}.$ ) . (2)
-We here briefly note that the probabilities $p_{n}$ are the eigenvalues of the density matrix $\rho_{\beta}=\frac{e^{-\beta H}}{\text{Tr}(e^{-\beta H})}$, where $H$ is the Hamiltonian of the (open) system.
-Using the density matrix $\rho_{\beta}$, the quantum Gibbs (qGB) entropy in Eq.2 can also be expressed in a basis-independent form, as discussed in Sec.III.1.
-It may be emphasized that the relevant density matrix is found here indirectly: starting from a classical (open) system, we first find the probability distribution and then reinterpreting the quantities in terms of the quantum variables, we get the density matrix.
-Instead, it is also possible to get the density matrix $\rho_{\beta}$ directly, starting with a quantum system [10, 11].
-In this work, we prefer the first method since it directly reveals when and how the TH entropy is equivalent to the qGB entropy.
-III von Neumann Entropy
+Subsequently, we can write the quantum version of the GB entropy (qGB entropy) in the following way:
+$S_{GB}\to S_{qGB}=-\sum_{n}p_{n}\ln{p_{n}}.$ (2)
+We here briefly note that the probabilities $p_{n}$ are the eigenvalues of the density matrix $\rho_{\beta}=\frac{e^{-\beta H}}{\mathrm{Tr}(e^{-\beta H})}$, where $H$ is the Hamiltonian of the (open) system.
+Using the density matrix $\rho_{\beta}$, the quantum Gibbs (qGB) entropy in Eq.2 can also be expressed in a basis-independent form, which looks the same as the VN entropy formula as appears in Eq.
+3.
+It may be emphasized that the quantum statistical entropy (qBN or qGB) of an isolated or open system can be put into a basis-independent form, appearing exactly the same as the VN entropy formula [17].
+But to show the equivalence between the VN entropy and the TH entropy (represented by qBN or qGB entropy), one should evaluate them independently.
+In Sec.
+4, we calculate them separately across all three cases mentioned earlier.
+3 von Neumann Entropy
 While reformulating some of the major aspects of statistical mechanics so that it can be used in studying quantum systems, von Neumann provided a formula for entropy of a quantum system in a mixed state [36].
 If the density matrix $\rho$ represents the mixed state of a system (not to be confused with the probability distribution function denoted by the same symbol), the entropy of the system, according to von Neumann, is
-$S_{VN}=-k_{B}\text{Tr}(\rho\ln\rho).$ (3)
+$S_{VN}=-\mathrm{Tr}(\rho\ln\rho).$ (3)
 This formula attracted both the applause and criticism alike [31, 13, 14, 7, 35].
-In the following subsections, we first discuss two different perspectives to understand how one can arrive at the formula and what it actually quantifies.
+In the following subsections, we first discuss two different perspectives to understand how one can arrive at this formula and what it actually quantifies.
 Next we highlight some of the major criticisms on the applicability of the formula and explain our views that many of such criticisms can be avoided or can be addressed convincingly.
-III.1 As basis-independent representation of quantum statistical entropy
-One way to understand the VN entropy is to view it as the basis-independent representation of the quantum statistical entropy of a thermodynamic system in equilibrium.
-The VN entropy for an appropriate density matrix can give us the qBN or qGB entropy.
-In fact, while deriving the expression of the thermodynamic (TH) entropy of a subsystem, Landau also came up with the same expression that appears in Eq.
-3 as the basis-independent representation of the quantum statistical entropy [9].
-For an isolated system, described by a microcanonical (mc) ensemble, we get the following from Eq.
-1: $S_{qBN}=-k_{B}\ln D=-k_{B}\text{Tr}(\rho_{mc}\ln\rho_{mc})=S_{VN}$, where $\rho_{mc}=\frac{1}{D}I$ with $D$ being the dimension of the relevant subspace of the Hilbert space of the system and $I$ is the identity matrix.
-So we find here that the VN entropy can be viewed as a basis-independent representation of the qBN entropy of an isolated system in thermal equilibrium.
-Similarly, for an open system, we get the following from the Eq.
-2, $S_{qGB}=-k_{B}\sum_{n}p_{n}\ln{p_{n}}=-k_{B}\text{Tr}(\rho_{\beta}\ln\rho_{\beta})=S_{VN}$, where $\rho_{\beta}=\frac{e^{-\beta H}}{\text{Tr}(e^{-\beta H})}$.
-Here $H$ and $\beta$ are, respectively, the Hamiltonian and the inverse temperature of the open system.
-We note again that the VN entropy of an open system in equilibrium can be viewed as a basis-independent representation of the qGB entropy.
-III.2 As entropy of thermodynamic ensemble
+3.1 Entropy of thermodynamic ensemble
 Sometimes we may not know in which pure state our quantum system is in; von Neumann introduced the concept of the density matrix ($\rho$) to describe the physics of such system [36].
@@
 As in Sec.
-II.3, if a pure quantum state is taken to be a microstate, then the qBN entropy of a system in a pure quantum state is zero (see Eq.
+2.3, if a pure quantum state is taken to be a microstate, then the qBN entropy of a system in a pure quantum state is zero (see Eq.
 1).
@@
 If we now see the decomposition in Eq.
-4, it would suggest that the corresponding entropy of the system is $S_{SN}(\{p_{i}\})=-k_{B}\sum_{i}p_{i}\ln p_{i}$, which is the Shannon (SN) entropy for the probability distribution $\{p_{i}\}$.
+4, it would suggest that the corresponding entropy of the system is $S_{SN}(\{p_{i}\})=-\sum_{i}p_{i}\ln p_{i}$, which is the Shannon (SN) entropy for the probability distribution $\{p_{i}\}$.
 Unfortunately, the decomposition of a mixed state is not unique, hence this way of defining entropy is unacceptable.
@@
 If we take a set of orthonormal states to be a set of distinguishable microstates, we can define an entropy from the probabilities associated with these microstates.
-If {|$\{\ket{\phi_{n}}\}$ is a set of orthonormal states, the entropy one can define is $S_{SN}(\{w_{i}\})=-k_{B}\sum_{n}w_{n}\ln w_{n}$, where $w_{n}=\bra{\phi_{n}}\rho\ket{\phi_{n}}$.
+If {|$\{\ket{\phi_{n}}\}$ is a set of orthonormal states, the entropy one can define is $S_{SN}(\{w_{i}\})=-\sum_{n}w_{n}\ln w_{n}$, where $w_{n}=\bra{\phi_{n}}\rho\ket{\phi_{n}}$.
 Unfortunately, again, this entropy is not unique as it depends on the set of orthonormal states (microstates).
@@
 5.
-For this updated density matrix after measurement ($\rho\to\rho^{\prime}$), as mentioned earlier, von Neumann showed that $-\text{Tr}(\rho^{\prime}\ln\rho^{\prime})\geq-\text{Tr}(\rho\ln\rho)$ [8].
+For this updated density matrix after measurement ($\rho\to\rho^{\prime}$), as mentioned earlier, von Neumann showed that $-\mathrm{Tr}(\rho^{\prime}\ln\rho^{\prime})\geq-\mathrm{Tr}(\rho\ln\rho)$ [36].
 Thus we see that the VN entropy corresponds to the minimum of $S_{SN}(\{w_{i}\})$, which we obtain when we choose the microstates as the eigenkets of the given density matrix $\rho$.
@@
 where it is understood that $\bra{\phi_{i}}\ket{\phi_{j}}=\delta_{ij}$.
-III.3 Addressing objections against VN entropy
+3.2 Addressing objections against VN entropy
 Over the century since von Neumann’s work, the entropy after his name found a lot of applications.
@@
 It is known that, for an isolated system, the TH entropy increases as the system approaches thermal equilibrium.
-Does not the VN entropy apply to systems in non-equilibrium?
+Does not the VN entropy apply to quantum systems in non-equilibrium?
 At the outset, we stress that the above issue arises only when the system is in a non-equilibrium state and we want to know the entropy during a process.
 In the thermal equilibrium, as discussed earlier in Sec.
-III.1, the VN entropy and the TH entropy are equivalent (under certain general conditions).
+2 and analyzed in details in Sec.
+4, the VN entropy and the TH entropy are equivalent (under certain general conditions).
 So, the VN entropy is capable of determining the change in (TH) entropy due to a process (the system is in equilibrium before and after the process).
@@
 The observational entropy is one such popular alternative to the VN entropy [35, 37, 21].
-Second, instead of considering that the VN entropy formula is wrong, one may consider that the time evaluation of the density matrix $\rho$ is not correctly described by the von Neumann-Liouville equation.
+Second, instead of considering that the VN entropy formula needs modification, one may consider that the time evaluation of the density matrix $\rho$ is not correctly described by the von Neumann-Liouville equation.
 We prefer the second option for the following reasons.
@@
 The second major objection against the VN entropy is that it obeys the subadditivity property while the TH entropy is additive (for a large uniform system where the interaction energy between its two parts can be ignored).
-The subadditivity property is stated as follows: if a bipartite system ($AB$) is in state $\rho_{AB}$, and its parts $A$ and $B$ are respectively in states $\rho_{A}$ and $\rho_{B}$ (which are reduced density matrices of the state $\rho_{AB}$), then $S_{VN}(\rho_{AB})\leq S_{VN}(\rho_{A})+S_{VN}(\rho_{B})$ [26].
+The subadditivity property is stated as follows: if a bipartite system ($AB$) is in state $\rho_{AB}$, and its parts $A$ and $B$ are respectively in states $\rho_{A}$ and $\rho_{B}$ (which are reduced density matrices from the state $\rho_{AB}$), then $S_{VN}(\rho_{AB})\leq S_{VN}(\rho_{A})+S_{VN}(\rho_{B})$ [2].
 This implies that, if the system is in a pure state, its VN entropy will be zero but the sum of the VN entropies of its parts is positive (the sum can be zero in a very special case).
@@
 It is known that the VN entropy follows a volume law for excited states (especially the states from the middle of its spectrum) [18, 1, 20, 3].
-In fact, for our system with $\Delta_{2}=0.5$ (see Sec.
-IV.3), we see that the average VN entropy $\overline{S}^{sb}_{VN}$ (averaged over all the states from a narrow energy window in the middle of the spectrum) is extensive and grows linearly with subsystem size; relevant results can be found in Fig.
+In fact, for our system with $\Delta_{2}=0.5$ (Sec.
+4.4), we see that the average VN entropy $\overline{S}^{sb}_{VN}$ (averaged over all the states from a narrow energy window in the middle of the spectrum) is extensive and grows linearly with subsystem size; relevant results can be found in Fig.
 1.
-Figure 1: In the middle of the spectrum, the average VN entropy ($\overline{S}^{sb}_{VN}$) grows linearly with subsystem size. Here total system size is $N=16$ and $\Delta_{2}=0.5$.
+Figure 1: In the middle of the spectrum, the average VN entropy ($\overline{S}^{sb}_{VN}$) grows linearly with subsystem size.
+The total system is described by the Hamiltonian of Eq.
+20 and $\overline{S}^{sb}_{VN}$ is calculated as per Eq.
+4.4.
+The calculations are done with the total system size $N=16$ and $\Delta_{2}=0.5$.
 We, therefore, infer here that the VN entropy also obeys the additivity property if interpreted appropriately.
@@
 Peres when he says that there should be no doubt that von Neumann’s entropy is equivalent to the entropy of classical thermodynamics [24].
-IV Entropy of large quantum systems
-In Sec.
-III.1, we have seen that the VN entropy of a thermalized isolated or open system is a basis-independent representation of the quantum statistical entropy (which, under certain conditions, is equivalent to the TH entropy).
-There we start with the qBN or qGB entropy and then show that it is simply equal to the VN entropy corresponding to some density matrix (we were not much concerned about any significance or possible interpretation of the density matrix).
-In this section, we go in the opposite direction: we first calculate the VN entropy independently and then verify if that is equal (equivalent) to the TH entropy of the system (subsystem).
+4 Entropy of large quantum systems
+In this section, we separately calculate the VN entropy and the TH entropy (represented by qBN or qGB entropy) of the system (or subsystem), and compare if they are equal or equivalent.
 To calculate the VN entropy, we take an appropriate density matrix which describes the system (subsystem) in thermodynamic equilibrium.
-In the following, we discuss the issue of equivalence separately for the isolated, open and subsystem (of a much larger system).
-IV.1 Isolated quantum system
+In the following, we discuss the issue of equivalence separately for the isolated, open, and subsystem.
+4.1 Isolated quantum system
 We consider an isolated quantum system in thermal equilibrium.
 As discussed in Sec.
-II.3, a large isolated quantum system will be in a mixed state, for one or multiple reasons.
+2.3, a large isolated quantum system will be in a mixed state, for one or multiple reasons.
 An isolated system in thermal equilibrium is appropriately described by a microcanonical (mc) ensemble.
-Assume that we prepare the system in such a way that its energy lies in the narrow energy shell $(E,E+\Delta E]$, where $\Delta E$ is arbitrary but macroscopically small and microscopically large [5].
+Assume that we prepare the system in such a way that its energy lies in the narrow energy shell $(E-\frac{\Delta E}{2},E+\frac{\Delta E}{2}]$, where $\Delta E$ is arbitrary but macroscopically small and microscopically large [19].
 Let {|$\{\ket{n}\}$ be the set of eigenkets of the Hamiltonian of the system, and $E_{n}$ be the eigenvalue corresponding to |$\ket{n}$.
 The density matrix which describes the isolated system is given by,
-$\rho_{mc}=\sum_{\begin{subarray}{c}n\\ E_{n}\in\left(E,E+\Delta E\right]\end{subarray}}p_{n}\ket{n}\bra{n},$ (8)
-where $p_{n}$ is the probability to find the system in state |$\ket{n}$.
+$\rho_{mc}={\sum_{n}}^{\prime}p_{n}\ket{n}\bra{n},$ (8)
+where $p_{n}$ is the probability to find the system in state |$\ket{n}$ and the primed sum is for the states associated with the sell $(E-\frac{\Delta E}{2},E+\frac{\Delta E}{2}]$.
 Since $\Delta E$ is macroscopically small, all the accessible microstates (here, the eigenkets) will essentially have almost the same energy.
-Application of the principle of equal a priori probabilities suggests that $p_{n}=\frac{1}{D}$ (see aslo Sec.
-II.3).
-As $S_{VN}=-k_{B}\text{Tr}(\rho_{mc}\ln\rho_{mc})$, we need the eigenvalues of $\rho_{mc}$ to efficiently compute $S_{VN}$.
+Application of the principle of equal a priori probabilities suggests that $p_{n}=\frac{1}{d_{E}}$ (see aslo Sec.
+2.3).
+The VN entropy of the system is now given by $S_{VN}=-\mathrm{Tr}(\rho_{mc}\ln\rho_{mc})$.
+One needs the eigenvalues of $\rho_{mc}$ to efficiently compute $S_{VN}$.
 It is clear from Eq.
-8 that |$\ket{n}$ is an eigenket of $\rho_{mc}$ with eigenvalue $\frac{1}{D}$, i.e., $\rho_{mc}\ket{n}=\frac{1}{D}\ket{n}$ for any energy eigenket from the shell.
+8 that any energy eigenket |$\ket{n}$ is an eigenket of $\rho_{mc}$ with eigenvalue $\frac{1}{d_{E}}$, i.e., $\rho_{mc}\ket{n}=\frac{1}{d_{E}}\ket{n}$.
 This implies that
-$S_{VN}=-k_{B}\text{Tr}(\rho_{mc}\ln\rho_{mc})=k_{B}\ln D.$ (9)
-If we now directly calculate the TH entropy (or equivalently, the qBN entropy), following our discussion in Sec. II.3, it would be
-$S_{TH}\equiv S_{qBN}=k_{B}\ln D,$ (10)
-where $D$ is the number of eigenvalues within the shell.
-It may be noted here that one can subdivide the energy shell into different macrostates (subspaces), with one of them having the largest number of microststes (the largest dimension) [7, 5].
-For a thermodynamically large system, the dimension of this largest subspace, which corresponds to the thermal equilibrium, is essentially equal to the dimension of the shell ($D$).
-We see from Eqs.
+$S_{VN}=-k_{B}\mathrm{Tr}(\rho_{mc}\ln\rho_{mc})=\ln d_{E}.$ (9)
+If we now directly calculate the TH entropy (or equivalently, the qBN entropy), following our discussion in Sec. 2.3, it would be
+$S_{TH}\equiv S_{qBN}=\ln d_{E},$ (10)
+where $d_{E}$ is the number of eigenvalues within the shell.
+Although $S_{VN}$ and $S_{TH}$ are here calculated separately, we see from Eqs.
 9 and 10 that, for an isolated quantum system, the VN entropy is the same as the TH entropy ($S_{VN}=S_{qBN}\equiv S_{TH}$), provided that the system is large and in thermal equilibrium.
-The detailed discussion on the conditions under which this equivalence works can be found in Sec.
-II.2.
-IV.2 Weakly interacting open quantum system
+The detailed discussion on the conditions under which this equivalence works can be found in Secs.
+2.2 and 2.3.
+4.2 Weakly interacting open quantum system
 Here we consider an open system that is in contact with the thermal bath.
 Although the system exchanges heat with the bath, we assume that the interaction with the bath is so weak that the energy spectrum of the system is more or less unaffected by the interaction, and the dynamics of the quantum system can still be described (at-least effectively) by its Hamiltonian $H$.
-As discussed in Sec.
-II.3, the density matrix that appropriately describes the open system is $\rho_{\beta}=\frac{e^{-\beta H}}{Tr(e^{-\beta H})}$, where the inverse temperature can be fixed by solving the equation $Tr(He^{-\beta H})=U$ with $U$ being the (given) internal or average energy.
-In fact, if the system-bath combination is described by a microcanonical ensemble or for the overwhelmingly large number of pure states (of the system-bath combination), it can be quite generally shown that an open quantum system is very well described by a (reduced) density matrix which is equivalent to the known Gibbsian distribution [11, 10].
+If the system-bath combination is described by a microcanonical ensemble or for the overwhelmingly large number of pure states (of the system-bath combination), it can be quite generally shown that an open quantum system is very well described by a (reduced) density matrix which is equivalent to the known Gibbsian distribution [26, 11].
+Denoting the density matrix by $\rho_{\beta}$, we have $\rho_{\beta}=\frac{e^{-\beta H}}{Tr(e^{-\beta H})}$, where the inverse temperature can be fixed by solving the equation $Tr(He^{-\beta H})=U$ with $U$ being the (given) internal or average energy.
 Accordingly, the VN entropy of the open system is,
-$S_{VN}=-k_{B}\text{Tr}(\rho_{\beta}\ln\rho_{\beta})=-k_{B}\sum_{n}p_{n}\ln{p_{n}},$ (11)
+$S_{VN}=-\mathrm{Tr}(\rho_{\beta}\ln\rho_{\beta})=-\sum_{n}p_{n}\ln{p_{n}},$ (11)
 where $p_{n}=\frac{1}{Q}e^{-\beta E_{n}}$ with $Q$ being the partition function, $Q=\sum_{n}e^{-\beta E_{n}}$. Here, $\{E_{n}\}$ is the set of eigenvalues of the system Hamiltonian $H$.
-Figure 2: The subsystem VN entropy ($S^{sb}_{VN}$) for the individual eigenkets and $\ln(DOS)$ are plotted against the energy of the full system. Here total system size is 16 and subsystem size is 6.
-As discussed in Sec. II, for the open quantum system, the TH entropy (or equivalently, the qGB entropy) is given by:
-$S_{TH}\equiv S_{qGB}=-k_{B}\sum_{n}p_{n}\ln{p_{n}},$ (12)
-where the expression of $p_{n}$ is the same as the one appears in Eq. 11.
-Figure 3: The average VN entropy ($\overline{S}^{sb}_{VN}$) of subsystem is plotted against $\ln(DOS)$ across the energy spectrum (separately for the left and the right halves of the spectrum).
-Here, the system (subsystem) size is 16 (6) and $\Delta_{2}=0.5$.
+As discussed in Sec. 2, for the open quantum system, the TH entropy (or equivalently, the qGB entropy) is given by:
+$S_{TH}\equiv S_{qGB}=-\sum_{n}p_{n}\ln{p_{n}},$ (12)
+where the expression of $p_{n}$ is the same as for the one appears in Eq. 11.
 We now see from Eqs.
 11 and 12 that the VN entropy is the same as the TH entropy ($S_{VN}=S_{qBN}\equiv S_{TH}$) for a weakly interacting open quantum system in thermal equilibrium.
-The detailed discussion on the conditions under which this equivalence works can be found in Sec.
-II.2.
-IV.3 Subsystem of a quantum system
-We separately discuss here the question of equivalence between the subsystem entropies when the full system is in a mixed state (described by a microcanonical ensemble) or in a pure state (both typical and atypical).
+The detailed discussion on the conditions under which this equivalence works can be found in Secs.
+2.2 and 2.3.
+4.3 Subsystem of a quantum system
+We now discuss the question of equivalence between the VN entropy and TH entropy of a subsystem.
+We separately discuss two different cases - when the full system is in a mixed state (described by a microcanonical ensemble) and when the full system is in a pure state.
+In the following, the full system is assumed to have two parts - the smaller part is called the “subsystem” (sb) and the rest of the system is called the “bath” (B).
 If the full system is in a mixed state $\rho_{mc}$, described by a microcanonical (mc) ensemble, then the subsystem will be in a Gibbsian distribution $\rho_{\beta}$ [26, 11].
-If $\rho_{sb}$ is the reduced density matrix for the subsystem, i.e., $\rho_{sb}=\text{Tr}_{B}(\rho_{mc})$, then $\rho_{sb}\approx\rho_{\beta}=\frac{e^{-\beta H}}{Q}$ with $Q=\text{Tr}_{sb}(e^{-\beta H})$.
-It may be noted that the first trace (Tr) operation here is over the degrees of freedom of the bath, B (complement of the subsystem in the full system), and the second trace operation is over the degrees of freedom of the subsystem (sb).
-We also note that, here, $H$ is the Hamiltonian of the subsystem if the interaction between the subsystem and the bath can be ignored, else it is the effective Hamiltonian of the subsystem, determined by $H=-\frac{1}{\beta}\ln(Q\text{Tr}_{B}(\rho_{mc}))$.
-This equation can be solved self-consistently: initially, we calculate $Q$ from the Hamiltonian of the subsystem (without interaction with the bath) and get the first order approximation of the effective $H$.
-Next we use this effective $H$ in $Q$ to find the next order approximation of the effective $H$.
-We continue this till we get a self-consistent result.
-Fortunately, we do not need to calculate this effective $H$ in the numerical calculations that follow.
-If the full system is in a pure state, for the overwhelming majority of pure states (of the total system), the subsystem can be described by a Gabssian distribution $\rho_{\beta}$ [11, 10].
-This important result, known as canonical typicality, leaves open the possibility of pure states (of the full system) for which a subsystem is not described by a Gibbsian distribution or for which a subsystem does not thermalize.
-We note that there are systems whose eigenstates (from a narrow energy window) are such that the reduced density matrices calculated from them are quite different to each other.
+If $\rho_{sb}$ is the reduced density matrix for the subsystem, i.e., $\rho_{sb}=\mathrm{Tr}_{B}(\rho_{mc})$, then $\rho_{sb}\approx\rho_{\beta}=\frac{e^{-\beta H}}{Q}$ with $Q=\mathrm{Tr}_{sb}(e^{-\beta H})$.
+It may be noted that the first trace (TrB) operation here is over the degrees of freedom of the bath (B), and the second trace operation (Trsb) is over the degrees of freedom of the subsystem (sb).
+We also note that, here, $H$ is the Hamiltonian of the subsystem and $\beta$ is the effective temperature determined by solving the following equation: $\lambda U=\mathrm{Tr}(H\rho_{\beta})$.
+Here $U$ is the given (fixed) internal energy of the full system and $\lambda$ is the ratio of the subsystem size to the full system size.
+Now following our discussion for the open system (see Sec.
+4.2), we can conclude that if the full system is described by a microcanonical ensemble, the VN entropy of a subsystem is equivalent to its TH entropy (i.e., qGB entropy).
+We also show this equivalence numerically in the next subsection (Sec.
+4.4).
+For this numerical verification, we independently calculate the VN entropy (as per Eq.
+4.4) and the TH entropy (as $\ln(DOS)$; see Eq.
+21) of the subsystem.
+We find that these two are linearly proportional to each other (as in Fig.
+3), confirming their equivalence (also see Sec.
+4.4).
+We now discuss when and how the equivalence between the VN entropy and the TH entropy of a subsystem holds if the full system is in a pure state.
+We take two different approaches here.
+In the first approach, we start with the landmark results of Page [22] and Sen [30].
+In the second approach, we appeal to the concept of canonical typicality [26, 11].
+If a quantum system of Hilbert space dimension $D_{1}D_{2}$ is in a random pure state, then the average VN entropy of the subsystem of Hilbert space dimension $D_{1}$ is
+$\mathbb{E}({S}^{sb}_{VN})=\left(\sum_{k=D_{2}+1}^{D_{1}D_{2}}\frac{1}{k}\right)-\frac{D_{1}-1}{2D_{2}}.$ (13)
+In the thermodynamic limit (with $1\ll D_{1}\ll D_{2}$), this result, first conjectured by Page [22] and later proven by Sen [30], becomes
+$\mathbb{E}({S}^{sb}_{VN})=\ln(D_{1})+O(D_{1}/D_{2}).$ (14)
+We argue here that, for a quantum system obeying ETH, the result of Eq.
+14 can be strengthened, and further argue that this stronger result implies equivalence of the VN entropy and the TH entropy for a subsystem.
+Consider the subspace spanned by the eigenkets of the total system corresponding to a narrow energy shell: $\mathcal{H}_{E}=\mathrm{Span}\{\ket{n}|E_{n}\in(E+\frac{\Delta E}{2},E-\frac{\Delta E}{2}]\}$, where $E_{n}$ is the eigenvalue corresponding to the eigenket |$\ket{n}$ of the total system.
+The shell width $\Delta E$, as mentioned earlier, is arbitrary, but is microscopically small and microscopically large.
+Let |$\ket{n_{1}},\ket{n_{2}}\in\mathcal{H}_{E}$, and corresponding reduced density matrices for the subsystem are $\rho_{1}=\mathrm{Tr}_{B}(\ket{n_{1}}\bra{n_{1}})$ and $\rho_{2}=\mathrm{Tr}_{B}(\ket{n_{2}}\bra{n_{2}})$.
+According to ETH, $\mathrm{Tr}_{sb}(\rho_{1}A)\approx\mathrm{Tr}_{sb}(\rho_{2}A)$, where $A$ is any local operator associated with the subsystem (sb).
+This hypothesis implies that the sets of eigenvalues of $\rho_{1}$ and $\rho_{2}$ can not be much different, and consequently, $S^{sb}_{VN}(\rho_{1})\approx S^{sb}_{VN}(\rho_{2})$, where $S^{sb}_{VN}(\rho):=-\mathrm{Tr}_{sb}(\rho\ln\rho)$ is the VN entropy of the subsystem corresponding to the reduced density matrix $\rho$.
+This result is also evident from Fig.
+2.
+Figure 2: The subsystem VN entropy ($S^{sb}_{VN}$) for the individual eigenkets and $\ln(DOS)$ are plotted against the energy of the full system (described by the Hamiltonian of Eq.
+20).
+Here total system size is 16 and subsystem size is 6.
+When we randomly choose states (of the full system) without restriction on the energy scale $E$, then all the energy scales associated with the subsystem are important.
+Accordingly, all states ($D_{1}$ in number) of the subsystem would contribute to its average VN entropy.
+This is an intuitive way to understand the result in Eq.
+14.
+When attention is restricted to a pure states in the subspace $\mathcal{H}_{E}$ associated with a narrow energy shell, all the states of the subsystem are no more equally significant.
+If $l_{1}$ ($l_{2}$) is the size of the subsystem (bath), then the states of the subsystem corresponding to the energy scale $\frac{l_{1}}{l_{1}+l_{2}}E$ would be important.
+Let $\widetilde{d}_{1}$ is the effective number of states of the subsystem corresponding to its energy scale $\frac{l_{1}}{l_{1}+l_{2}}E$.
+We can also view $\widetilde{d}_{1}$ as the dimension of the subsystem’s effective Hilbert space associated with the subspace $\mathcal{H}_{E}$.
+We then expect:
+$\overline{S}^{sb}_{VN}\approx\ln\widetilde{d}_{1},$ (15)
+where the average is over all eigenkets from the subspace $\mathcal{H}_{E}$. To estimate $\widetilde{d}_{1}$, we note that
+$\widetilde{d}_{1}\widetilde{d}_{2}=d_{E},$ (16)
+where $\widetilde{d}_{2}$ is similarly defined dimension of the effective Hilbert space of the bath (B) associated with the subspace $\mathcal{H}_{E}$ and $d_{E}$ is the dimension of the subspace, i.e., $d_{E}:=\dim{(\mathcal{H}_{E})}$.
+We now define a parameter $\eta$ (may be called the effective Hilbert space dimension of a single spin) so that $\eta^{l_{1}}:=\widetilde{d}_{1}$ and $\eta^{l_{2}}:=\widetilde{d}_{2}$.
+This is similar to $d^{l_{1}}=D_{1}$ and $d^{l_{2}}=D_{2}$ where $d$ is the Hilber space dimension of a single spin (for spin-1/2 system, $d=2$).
+This shows that
+$\frac{\ln\widetilde{d}_{1}}{\ln\widetilde{d}_{2}}=\frac{\ln D_{1}}{\ln D_{2}}.$ (17)
+Now starting from Eq.
+16, and using Eq.
+17, we get, $\ln d_{E}=\ln\widetilde{d}_{1}(1+\frac{\ln\widetilde{d}_{2}}{\ln\widetilde{d}_{2}})=\ln\widetilde{d}_{1}(1+\frac{\ln D_{2}}{\ln D_{1}})=\ln\widetilde{d}_{1}\frac{\ln D}{\ln D_{1}}$.
+Rearranging the terms, we find, $\ln\widetilde{d}_{1}=\ln D_{1}(\frac{\ln d_{E}}{\ln D})$.
+Finally,
+$\widetilde{d}_{1}=D_{1}^{\gamma},\mathrm{~{}with~{}}\gamma=\frac{\ln d_{E}}{\ln D}.$ (18)
+This finding confirms our proposition that the result in Eq.
+14 can be strengthened for a quantum system that obeys ETH.
+In other words, for a quantum system which thermalizes, the result of Page and Sen holds for the subspaces associated with each narrow energy window.
+We may here naively note that the limit $\gamma=1$ gives the results of Page and Sen.
+Figure 3: The average VN entropy ($\overline{S}^{sb}_{VN}$) of subsystem is plotted against $\ln(DOS)$ across the energy spectrum (separately for the left and the right halves of the spectrum).
+Here the total system is described by the Hamiltonian of Eq.
+20 and $\overline{S}^{sb}_{VN}$ is calculated as per Eq.
+4.4.
+The calculations are done with the system (subsystem) size 16 (6) and $\Delta_{2}=0.5$.
+To get numerical evidence to assert our proposition (Eq. 15), we first note that
+$\overline{S}^{sb}_{VN}\approx\ln\widetilde{d}_{1}=(\frac{\ln D_{1}}{\ln D})\ln d_{E}=\lambda\ln d_{E},$ (19)
+where $\lambda$ is the ratio of the size of the subsystem to the size of the total system (i.e., $\frac{l_{1}}{l_{1}+l_{2}}$).
+If we take unit energy interval, we find from Eq.
+19 that $\overline{S}^{sb}_{VN}\propto\ln(DOS)$, where $DOS$ is the density-of-states of the full system at energy $E$.
+For numerical calculations, we take a spin-1/2 one-dimensional system with next-nearest neighbor interactions, as described by the Hamiltonian in Eq.
+20.
+For each narrow energy interval, we evaluate $\overline{S}^{sb}_{VN}$ and $\ln(DOS)$ independently and plot them for different energies across the energy spectrum of the full system, as seen in Fig.
+3.
+The linearity of the plot asserts our proposition that, for a system which obeys ETH, the result of Page and Sen holds for the subspaces associated with each narrow energy window.
+Interestingly, the result presented above can also be interpreted as the statement of equivalence between the VN entropy and the TH entropy of a subsystem (of a larger system which thermalizes).
+We note that $\ln d_{E}$ is the TH entropy (more specifically, qBN entropy) of the full system at energy $E$.
+Assuming the additivity of the TH entropy, one can say that the TH entropy of the subsystem is proportional to $\ln d_{E}$.
+This finally shows, from Eq.
+19, that the VN entropy and the TH entropy of a subsystem are equivalent.
+We now discuss the issue of equivalence using the concept of the canonical typicality.
+If the full system is in a pure state, for the overwhelming majority of pure states (of the total system), the subsystem can be described by a Gibbsian distribution $\rho_{\beta}$ [26, 11].
+This important result immediately implies that, if the full system is in a pure state, for the overwhelming cases (i.e., for the typical states), the VN entropy of a subsystem will be equivalent to its TH entropy (qGB entropy), as in the case of an open system.
+The result of the canonical typicality leaves open the possibility of pure states (of the full system) for which a subsystem is not described by a Gibbsian distribution or for which a subsystem does not thermalize.
+We note that there are systems whose eigenstates (from a narrow energy window) are such that the eigen spectra of the corresponding reduced density matrices are quite different to each other.
+The VN entropy calculated from them has large deviation from the mean (see Fig.
+2).
 These systems do not follow the eigenstate thermalization hypothesis or ETH [5, 33, 19].
-Above important results immediately imply the following: if the full system is described by a microcanonical ensemble, the VN entropy of a subsystem is equivalent to its TH entropy (qGB entropy).
-If the full system is in a pure state, for the overwhelming cases (i.e.
-for the typical states), the VN entropy of a subsystem will be equivalent to its TH entropy (qGB entropy).
-But for systems for which the ETH does not hold, this equivalence will not work in general (corresponding pure states are atypical).
-In the following we take a spin system and numerically demonstrate the issue of equivalence between the VN entropy and TH entropy of a subsystem in the different cases mentioned above.
-For our numerical results, we take the following one-dimensional (with open boundary condition) Heisenberg spin-1/2 chain with the next-nearest neighbor interactions:
+For such systems, the VN entropy and the TH entropy (or qBN entropy) are not equivalent.
+Figure 4: The subsystem VN entropy ($S^{sb}_{VN}$) when the full system is in a mixed state (see Eq. 4.4). Here, the system (subsystem) size is 16 (6).
+4.4 Numerical results
+In the part we take a spin system and numerically demonstrate the issue of equivalence between the VN entropy and the TH entropy of a subsystem in the different cases discussed earlier.
+For all our numerical results, we take the following one-dimensional (with open boundary condition) Heisenberg spin-1/2 chain with the next-nearest neighbor interactions:
 $H=\sum_{i=1}^{N-1}\vec{S}_{i}\cdot\vec{S}_{i+1}+\Delta_{2}\sum_{i=1}^{N-2}S^{z}_{i}S^{z}_{i+2}.$ (20)
 In our calculations, we take the size of the system $N=16$. This model system is integrable when $\Delta_{2}=0$ and is non-integrable (hence thermalizes) when $\Delta_{2}>0$ [34].
-To numerically calculate the VN entropy of the subsystem, we first find $\rho_{sb}$ from the state ($\rho$) of the full system: $\rho_{sb}=\text{Tr}_{B}(\rho)$.
-We then use the formula Eq.
-3 to calculate the VN entropy of the subsystem.
-To numerically calculate the corresponding TH entropy, consider that the full system is prepared with an energy lying in $(E,E+\Delta E]$.
-As discussed earlier, the system will be in a mixed state $\rho_{mc}$, as appear in Eq.
+To numerically calculate the TH entropy, consider that the full system is prepared with an energy lying in $(E-\frac{\Delta E}{2},E+\frac{\Delta E}{2}]$.
+The TH (or qBN) entropy of the full system is $S_{TH}\equiv S_{qBN}=\ln d_{E}$, as in Eq.
+10, where $d_{E}$ is the number of eigenvalues of the system lying in the aforementioned energy shell.
+Now since the TH entropy is additive, we can assume that the TH entropy of the subsystem (denoted as $S^{sb}_{TH}$ ) is proportion to $\ln d_{E}$ or to the $\ln(DOS)$, where $DOS$ is the density-of-states of the full system at the energy scale $E$.
+Accordingly,
+$S^{sb}_{TH}\equiv S^{sb}_{qBN}\propto\ln(DOS),$ (21)
+where the proportionality constant depends on the size of the subsystem in comparison to the size of the full system.
+Now while numerically calculating the VN entropy of the subsystem, we consider two separate cases - full system in a mixed state and full system in a pure state.
+First consider that the full system is in mixed state $\rho_{mc}$, as appear in Eq.
 8.
-The TH (or qBN) entropy of the full system is $S_{TH}\equiv S_{qBN}=k_{B}\ln D$, as in Eq.
-10, where $D$ is the number of eigenvalues of the system lying in the aforementioned energy shell.
-Now since the TH entropy is additive, we can assume that the TH entropy of the subsystem (denoted as $S^{sb}_{TH}$ ) is proportion to $k_{B}\ln D$ or to the $k_{B}\ln(DOS)$, where $DOS$ is the density-of-states of the full system at the energy scale $E$.
-Accordingly,
-$S^{sb}_{TH}\equiv S^{sb}_{qBN}\propto k_{B}\ln(DOS),$ (14)
-where the proportionality constant depends on the size of the subsystem in comparison to the size of the full system.
-Now, for different cases, we calculate the VN entropy and the TH entropy of a subsystem to study their equivalence.
-First, we consider the case when the full system is in an eigenket |$\ket{n}$ of the Hamiltonian of Eq.
-13.
-In this case the subsystem will be represented by the reduced density matrix $\rho_{sb}\equiv\rho_{sb}^{(n)}=\text{Tr}_{B}(\ket{n}\bra{n})$.
-Now using Eq.
-3, we can calculate the VN entropy of the subsystem when the full system is in the eigenket |$\ket{n}$:
-$S^{sb}_{VN}=-k_{B}\text{Tr}(\rho_{sb}\ln\rho_{sb}).$ (15)
-The values of $S^{sb}_{VN}$ for different eigenkets across the spectrum can be found in Fig.
+Let $\rho_{sb}$ be the corresponding reduced density matrix of the subsystem: $\rho_{sb}=\mathrm{Tr}_{B}(\rho_{mc})$.
+More specifically, $\rho_{sb}=\frac{1}{d_{E}}\sum^{\prime}_{n}\rho_{sb}^{(n)}$, where the reduced density matrix $\rho_{sb}^{(n)}=\mathrm{Tr}_{B}(\ket{n}\bra{n})$ is calculated for a single eigenket |$\ket{n}$ of the full system.
+Here the primed sum is over the all eigenkets associated with a small (fixed) energy shell $(E-\frac{\Delta E}{2},E+\frac{\Delta E}{2}]$.
+There are $d_{E}$ eigenkets associated with the shell.
+The VN entropy of the subsystem, then, would be $S^{sb}_{VN}=-\mathrm{Tr}_{sb}(\rho_{sb}\ln\rho_{sb})$.
+Thus, in the current scenario, we have the following:
+$\rho_{sb}$ $=$ $\frac{1}{d_{E}}{\sum_{n}}^{\prime}\rho_{sb}^{(n)},~{}\mathrm{with}~{}\rho_{sb}^{(n)}=\mathrm{Tr}_{B}(\ket{n}\bra{n})$
+$S^{sb}_{VN}$ $=$ $-\mathrm{Tr}_{sb}(\rho_{sb}\ln\rho_{sb}).$ (22)
+The plot of $S^{sb}_{VN}$ across the energy spectrum of the full system can be found in Fig.
+4.
+The plot of $S^{sb}_{VN}$ vs.
+$\ln(DOS)$ at different energy scale ($E$) is linear and very similar to the plot of $\overline{S}^{sb}_{VN}$ vs.
+$\ln(DOS)$ for the thermalized system as discussed below (and the corresponding plot can be found in Fig.
+3).
+The linearity confirms that the VN entropy and the TH entropy (or qBN entropy) of a subsystem are equivalent when the full system is in a mixed state (described by a microcanonical ensemble).
+Next, we consider the case where the full system is in a pure state.
+Once again, when the full system is described by a pure-state density matrix $\rho$, the reduced density matrix of the subsystem would be: $\rho_{sb}=\mathrm{Tr}_{B}(\rho)$.
+Consequently, the VN entropy of the subsystem is $S^{sb}_{VN}=-\mathrm{Tr}_{sb}(\rho_{sb}\ln\rho_{sb})$.
+Here we calculate this VN entropy corresponding to each eigenket from the pre-specified energy shell, and then find the average VN entropy $\overline{S}^{sb}_{VN}$.
+In the current scenario, we have the following:
+$\rho^{(n)}_{sb}$ $=$ $\mathrm{Tr}_{B}(\ket{n}\bra{n})$
+$\overline{S}^{sb}_{VN}$ $=$ $-\frac{1}{d_{E}}{\sum_{n}}^{\prime}\mathrm{Tr}_{sb}(\rho^{(n)}_{sb}\ln{\rho^{(n)}_{sb}}),$ (23)
+where the primed sum is over all the eigenkets associated with the shell $(E-\frac{\Delta E}{2},E+\frac{\Delta E}{2}]$.
+The values of $S^{sb}_{VN}$ for the individual eigenkets across the spectrum can be found in Fig.
 2 (for $\Delta_{2}=0.0,0.2$ and $0.5$).
 We also show the corresponding logarithm of the density-of-states (DOS) for the full system.
-As mentioned earlier, when the system does not thermalize ($\Delta_{2}=0.0$), the individual reduced density matrices are quite different from each other (even within a narrow energy shell) and accordingly, $S^{sb}_{VN}$ takes quite different values as one goes from one eigenket to another within the energy shell.
-In this case the VN entropy of the subsystem is not, in general, proportional to its qBN entropy, i.e., to $\ln(DOS)$.
+As mentioned earlier, when the system does not thermalize ($\Delta_{2}=0.0$), the eigen spectra of individual reduced density matrices are quite different from each other (even within a narrow energy shell) and accordingly, $S^{sb}_{VN}$ takes quite different values as one goes from one eigenket to another within the energy shell.
+In this case the VN entropy of the subsystem is not, in general, proportional to its qBN entropy, i.e., to $\ln(DOS)$ with proper scaling factor.
 On the other hand, when the system thermalizes (or the ETH holds), $S^{sb}_{VN}$ takes very close values as one goes from one eigenket to another within an energy shell (see, for example, results for $\Delta_{2}=0.5$).
 In Fig.
-3, we plot the average VN entropy ($\overline{S}^{sb}_{VN}$) of the subsystem against the logarithm of DOS (individually for left and right halves of the spectrum).
-We take the average of $S^{sb}_{VN}$ over a small energy shell to get better numerical results.
-The linearity of the plots shows that the VN entropy of a subsystem is equivalent to its TH (or qBN) entropy, i.e., $S^{sb}_{VN}\approx\overline{S}^{sb}_{VN}\equiv S^{sb}_{qBN}$.
-Next, we consider the case when the full system is in a mixed state ($\rho_{mc}$), as appears Eq. 8 (with $p_{n}=\frac{1}{D}$). In this case, the subsystem will be in the following mixed state,
-$\rho_{sb}=\frac{1}{D}\sum_{n}\rho_{sb}^{(n)},$ (16)
-where the reduced density matrix $\rho_{sb}^{(n)}=\text{Tr}_{B}(\ket{n}\bra{n})$ is calculated for a single eigenket |$\ket{n}$ of the full system.
-Here the sum is over the all eigenkets from a small (fixed) energy shell.
-There are $D$ eigenkets in the shell.
-Using Eq.
-15, we now can calculate the VN entropy of the subsystem directly from $\rho_{sb}$ of Eq.
-16.
-The value of $S^{sb}_{VN}$ across the energy spectrum can be found in Fig.
+3, we plot the average VN entropy ($\overline{S}^{sb}_{VN}$ as in Eq.
+4.4) of the subsystem against the logarithm of DOS (separately for the left and right halves of the spectrum).
+The linearity of this plot is as per our prediction in the last subsection (Sec.
+4.3), and it confirms that the VN entropy and the TH entropy (or qBN entropy) of a subsystem are equivalent when the full system is in pure state (this holds when the full system thermalizes).
+We conclude this subsection with a few remarks.
+When a system is thermalized, the VN entropy of its subsystem, corresponding to any individual eigenket from a narrow energy shell, is quite close to its mean value.
+This is evident in Fig.2(a).
+For any thermalized or non-thermalized system, when we average the reduced density matrices associated with a narrow energy shell, the eigen spectrum of this resultant density matrix is expected to be very similar to the eigen spectrum of a reduced density matrix associated with a thermalized system.
+As a consequence, the results from Eq.
+4.4, irrespective of the type of system, will be very similar to the results from Eq.
+4.4 when the system is thermalized.
+This is evident if one compares Fig.
+2(a) and Fig.
 4.
-Figure 4: The subsystem VN entropy ($S^{sb}_{VN}$) when the full system is in the microcanonical density matrix $\rho_{mc}$. Here, the system (subsystem) size is 16 (6).
-The plot of $S^{sb}_{VN}$ against $\ln(DOS)$ shows linearity, similar to what we find in Fig.
-3.
-We note that, as expected, we get the said linear plot even when the system is in integrable regime ($\Delta_{2}=0.0$).
-It may be worth mentioning here that, the result that the subsystem’s average VN entropy is proportional to system’s $\ln(DOS)$, is earlier reported in [36] in a different context.
-We also note that, elsewhere [37], the TH entropy of a subsystem is defined as what we call here the VN entropy of a subsystem ($S^{sb}_{VN}$) when the full system is in $\rho_{mc}$.
+We also remark here that, in a different context [28], numerically it was previously observed that the subsystem’s average VN entropy is proportional to the system’s $\ln(DOS)$.
+We also note that, elsewhere [6], the TH entropy of a subsystem is defined as what we call here the VN entropy of a subsystem ($S^{sb}_{VN}$) when the full system is in a mixed state ($\rho_{mc}$).
 As discussed in this work, the TH entropy of a subsystem is appropriately calculated as the logarithm of $DOS$ (since, $S^{sb}_{TH}\equiv S^{sb}_{qBN}\propto\ln(DOS)$ under some general conditions).
-V Conclusion
+5 Conclusion
 In this work we analyze whether and when the von Neumann (VN) entropy is equal or equivalent to the thermodynamic (TH) entropy.
 In reviewing the major objections about the VN entropy, we explain that the concerns about the time-invariance and subadditivity properties can be addressed convincingly.
-We find that, when a large isolated system is in a mixed state (described by microcanonical ensemble), its VN entropy is equivalent to its TH entropy.
-When the system-bath combination is in a mixed state, or in a typical pure state, the VN entropy of the (open) system is equivalent to its TH entropy.
-This result also holds for a subsystem (of a much larger system) but the analytical demonstration is not straightforward.
-Through numerical analysis, we show that the VN entropy of a subsystem is equivalent to its TH (qBN) entropy when the full system (subsystem-bath combination) is described by a microcanonical ensemble or when the full system is in a typical pure state.
-In this work we also numerically demonstrate that, for the states that are atypical (or for which the ETH does not hold), the VN entropy and the TH entropy of a subsystem are not the same.
-Acknowledgements.
+We separately address the question of equivalence for isolated system, open system, and subsystem (both when the full system is in a mixed state or in a pure state).
+In this context, we show that the result of Page and Sen can be strengthened for a thermalized system, and further argued that this new result implies the equivalence of the VN entropy and the TH entropy of a subsystem when the full system is in a (random) pure state associated with a narrow energy shell.
+We provide numerical data to demonstrate or assert some of our key results.
+6 Acknowledgement
 SS thanks S. Aravinda for useful discussions.